### v1

# Traveling supersolid stripe patterns in spin-orbit-coupled Bose-Einstein condensates

Abstract
We consider a traveling supersolid stripe pattern in a spin-orbit-coupled Bose gas. This configuration is associated with an unequal occupation of the two single-particle energy minima, giving rise to a chemical potential difference that sets the fringe velocity. Unlike stationary stripes, the moving pattern is spin-polarized, with decreasing contrast as momentum increases, eventually leading to stripe melting and transition to the uniform plane-wave phase. The Bogoliubov spectrum of the moving stripes exhibits asymmetry under inversion of the excitation quasimomentum. At high momentum, we identify energetic and dynamical instabilities in the spin-phonon mode which transforms to the roton mode of the plane-wave phase as the stripe structure vanishes.
I Introduction
In recent years, the interest in supersolidity, where superfluidity coexists with a crystal-like structure, has grown significantly (see reviews in [1, 2, 3, 4, 5, 6]). First discussed in the 1950s [7, 8, 9] and further developed in subsequent decades [10, 11, 12, 13, 14, 16, 15, 17, 18], the supersolidity has only recently been observed in ultracold boson systems inside optical resonators [19], with spin-orbit coupling [20, 21, 22], and with dipolar interactions [23, 24, 25, 26, 27], as well as in exciton-polariton systems [28, 29]. These advances sparked studies on several aspects of supersolids, including their ability to sustain a dissipationless flow, i.e., a supercurrent. At low flow velocities, the superfluid fraction–bounded strictly below one due to crystalline order [13, 18]–is inferred from the system’s response to a translation [30, 31, 32, 33, 34, 35, 36] or rotation [39, 37, 40, 38, 41] constraint. However, accessing this quantity in experiments has proven challenging [42, 43], and only very recently a measurement has been performed [44]. At higher flow velocities, supercurrent-carrying supersolids can become unstable [45, 46, 32, 47], and in rotating configurations this can lead to the formation of vortices [17, 48, 49, 50], as confirmed in recent experiments [51, 52]. An anomalous Doppler effect has also been predicted [53]. Interestingly, although these results indicate that supersolids are (under certain conditions) able to sustain a supercurrent, frictionless motion of external objects remains impossible [17, 54].
A common method to generate a supercurrent in a Bose-Einstein condensate (BEC) is to impose a phase twist on its time-independent order parameter [55]. In supersolids, this leads to a configuration where the lattice associated with the normal component remains at rest, while the superfluid background flows at a velocity determined by the twist angle. An alternative approach involves considering a moving lattice pattern with velocity $\bm{v}$, making the time-dependent order parameter a function of $\bm{r}-\bm{v}t$. In Galilean-invariant systems, such as dipolar gases, these two descriptions are physically equivalent, being connected by a Galilean transformation. However, spin-orbit-coupled BECs lack Galilean invariance, a feature that profoundly alters their superfluid behavior even in the non-supersolid phases. Early studies have shown that in these phases the superfluid fraction is strictly less than unity even at zero temperature [56, 57, 58]. In addition, the critical velocity for supercurrent stability differs from the threshold for frictionless impurity motion [59, 60], and current-carrying configurations can exhibit both energetic and dynamical instabilities [61], unlike in standard condensates, where only energetic (Landau) instabilities are present [62, 63]. In the supersolid stripe phase, the combined lack of translational and Galilean invariance further reduces the superfluid density compared to the Leggett bound [58] and leads to peculiar stability conditions for current-carrying states [64].
The Raman lasers responsible for generating spin-orbit coupling play a crucial role in the construction of supercurrent states within the stripe phase. In the analysis of Ref. [64], the current is generated by imposing twisted boundary conditions on the condensate order parameter. This procedure effectively constrains the system’s kinetic momentum and describes a scenario in which the superfluid background flows relative to the Raman laser frame, while the density modulations remain stationary. This is consistent with the idea that the Raman lasers pin the normal component [56]. In the rest frame of the superfluid, both the fringes and the lasers appear to move at the same velocity. These configurations correspond to the two contrasting, but physically equivalent, pictures discussed earlier: background flow versus fringe motion.
In this work, we unveil a third scenario, the one in which the density modulations and the Raman lasers move relative to each other. This relative motion renders the configuration physically distinct from the two previously discussed cases. Here, the relevant conserved quantity is the canonical momentum, which, unlike the kinetic momentum, commutes with the spin-orbit Hamiltonian [61]. In contrast to earlier scenarios, these traveling stripe patterns do not allow a simple interpretation in terms of independent motion of the superfluid or normal components. Instead, they can be understood as a generalization of the notion of an imbalanced BEC mixture to the context of spin-orbit coupling, where the relevant degrees of freedom are dressed spin states rather than bare spin components. They represent intermediate structures that smoothly interpolate between the stationary stripe phase, i.e., a balanced mixture of two dressed spin states, and the uniform plane-wave phase corresponding to a fully polarized configuration in dressed spin. Within this framework, the fringe motion arises from a population imbalance between the two dressed spin components entering the condensate order parameter. The resulting chemical potential difference drives the density fringes to propagate at a constant velocity.
The motion of density fringes relative to stationary Raman lasers is well understood in the low-velocity limit, where it corresponds to the crystal Goldstone mode of the stripe phase and can be excited by releasing a weak spin perturbation [65, 66]. The goal of this work is to systematically investigate the properties of traveling stripe configurations beyond this regime. We show that at finite momentum the moving stripes develop asymmetric density profiles in the two spin components. Moreover, they display finite spin polarization as well as reduced contrast and higher energy compared to stationary stripes, effects that intensify with increasing momentum. Eventually, the fringes disappear, and the system undergoes a transition to a uniform plane-wave condensate. An analysis of the excitation spectrum of the moving supersolid reveals both energetic and dynamical instabilities emerging at sufficiently high momentum, involving the spin phonon branch. This mode becomes the mode with the roton minimum of the plane-wave phase [67].
The mechanism considered here for generating a chemical potential difference through a population imbalance between dressed spin states relies on the presence of antiferromagnetic spin-dependent interactions. The resulting fringe motion can thus be regarded as a purely interaction-driven effect. A chemical potential mismatch can also be induced by an effective Zeeman splitting generated via finite Raman detuning, even when the two dressed spin components in the order parameter have equal populations. Although this scenario is not explored in the present work, we note that such a method was successfully used in the experiment of Ref. [20] to observe moving stripes.
This paper is structured as follows. After briefly reviewing the model for a BEC with Raman-induced spin-orbit coupling (Sec. II), we outline the construction of traveling stripe solutions (Sec. III). These solutions are then derived and analyzed in detail (Sec. IV), with numerical results compared with perturbative analytical estimates. Section V focuses on the Bogoliubov spectrum and the instabilities of the moving stripes. Conclusions are presented in Sec. VI. Technical details of the perturbative approach and full expressions of relevant coefficients appearing in the formulas of the main text are given in Appendices A and B, respectively.
II The model
Let us consider a spin-$1/2$ BEC with spin-orbit coupling. Within the mean-field approximation, the state of the system is described by a two-component order parameter $\Psi(\bm{r},t)=(\Psi_{\uparrow}(\bm{r},t),\Psi_{\downarrow}(\bm{r},t))^{T}$, whose time evolution is governed by the time-dependent Gross-Pitaevskii equation:
$\mathrm{i}\hbar\partial_{t}\Psi=h_{\mathrm{SO}}\Psi+g_{dd}\left(\Psi^{\dagger}\Psi\right)\Psi+g_{ss}\left(\Psi^{\dagger}\sigma_{z}\Psi\right)\sigma_{z}\Psi\,.$ (1)
Here, $h_{\mathrm{SO}}$ is the single-particle Hamiltonian incorporating spin-orbit coupling, and $\sigma_{x,y,z}$ the standard Pauli matrices. The interaction parameters $g_{dd}=(g+g_{\uparrow\downarrow})/2$ and $g_{ss}=(g-g_{\uparrow\downarrow})/2$ represent the density-density and spin-spin coupling strengths, respectively. In writing Eq. (1), we have assumed equal intraspecies interaction strengths, $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}\equiv g$. The nonlinear coupling constants are related to the corresponding $s$-wave scattering lengths through the standard relation $g_{\sigma\sigma^{\prime}}=4\pi\hbar^{2}a_{\sigma\sigma^{\prime}}/m$ ($\sigma,\sigma^{\prime}=\uparrow,\downarrow$), where $m$ is the atomic mass.
In the case of one-dimensional Raman-induced spin-orbit coupling, which is the focus of this work, the single-particle Hamiltonian takes the form [68]
$h_{\mathrm{SO}}=\frac{\left(p_{x}-\hbar k_{R}\sigma_{z}\right)^{2}}{2m}+\frac{p_{y}^{2}+p_{z}^{2}}{2m}+\frac{\hbar\Omega_{R}}{2}\,\sigma_{x}+\frac{\hbar\delta_{R}}{2}\,\sigma_{z}\,,$ (2)
where $\bm{p}=-\mathrm{i}\hbar\nabla_{\bm{r}}$ is the canonical momentum operator. The spin-orbit coupling strength is proportional to the momentum imparted by the Raman lasers, which is $-2\hbar k_{R}\hat{\bm{e}}_{x}$, with $\hat{\bm{e}}_{x}$ the unit vector along the $x$ axis. The corresponding energy scale is $E_{R}=\hbar^{2}k_{R}^{2}/2m$. The parameter $\Omega_{R}$ quantifies the Raman coupling strength, while $\delta_{R}$ represents the Raman detuning, which we shall set to zero from now on.
It is worth noting that Eq. (3) can be recast in the variational form $\mathrm{i}\hbar\partial_{t}\Psi=\delta E/\delta\Psi^{\dagger}$, where $E$ is the total energy functional of the system:
$E=\int_{V}\mathrm{d}^{3}r\left(\Psi^{\dagger}h_{\mathrm{SO}}\Psi+\frac{g_{dd}}{2}n^{2}+\frac{g_{ss}}{2}s_{z}^{2}\right)\,.$ (3)
Here, $n=\Psi^{\dagger}\Psi$ is the total particle density and $s_{z}=\Psi^{\dagger}\sigma_{z}\Psi$ is the spin density along the $z$ axis. The integral extends over the volume $V$ containing the condensate.
The two-component Gross-Pitaevskii equation (1) conserves two important quantities in addition to the total energy (3). First, the invariance of the equation under global phase rotations of the order parameter $\Psi$ implies the conservation of the particle number
$N=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\Psi\,.$ (4)
This condition corresponds to the standard normalization of $\Psi$.
Second, due to the space translation symmetry of Eq. (1), the expectation value of the canonical momentum,
$\langle\bm{p}\rangle=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\bm{p}\Psi\,,$ (5)
is conserved. We note that $\bm{p}$ differs from the kinetic momentum $\bm{P}=\bm{p}-\hbar k_{R}\sigma_{z}\hat{\bm{e}}_{x}$, that is, the operator whose expectation value, $\langle\bm{P}\rangle=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\bm{P}\Psi$, gives the center-of-mass velocity of the system multiplied by the total mass $Nm$.
Due to the term $-\hbar k_{R}\sigma_{z}$ arising from the spin-orbit coupling, the $x$ component of the kinetic momentum does not commute with the single-particle Hamiltonian (2), unless the Raman coupling $\Omega_{R}$ vanishes. This implies the lack of Galilean invariance in spin-orbit-coupled BECs, which has significant implications for the superfluid behavior, as discussed in the introduction. In the context of this work, the key implication is that constructing states with a constant superflow is not the same as constructing states that propagate at a constant velocity. The former involves fixing the kinetic momentum $\langle\bm{P}\rangle$, as in the calculations of the superfluid fraction in [56, 57] and in the study of current-carrying supersolid configurations in [64]. In contrast, here we focus on moving supersolid patterns with fixed canonical momentum $\langle\bm{p}\rangle$. For clarity we will henceforth refer to this quantity simply as “momentum,” omitting the qualifier “canonical.”
The ground state properties of an interacting spin-orbit-coupled BEC can be understood by first examining the structure of the single-particle energy spectrum. Upon diagonalization of the Hamiltonian (2) at fixed momentum $\bm{p}$, two energy branches emerge. The lower branch exhibits either two degenerate minima at momenta ±$\pm k_{1}^{\mathrm{SP}}\hat{\bm{e}}_{x}=\pm k_{R}\sqrt{1-(\hbar\Omega_{R}/4E_{R})^{2}}\hat{\bm{e}}_{x}$ (for $\hbar\Omega_{R}<4E_{R}$) or a single minimum at zero momentum (when $\hbar\Omega_{R}\geq 4E_{R}$[68]. The interplay between this structure and interaction effects gives rise to a rich equilibrium phase diagram. The phase structure can be inferred by minimizing the total energy (3) as a function of the system parameters [69, 70] (for comprehensive overviews, see also the reviews in Refs. [71, 72, 73, 74, 75]). Among the three quantum phases that emerge, the stripe phase is of particular interest. This phase appears at low Raman coupling in the presence of an antiferromagnetic spin-dependent interaction, i.e., when $g_{ss}>0$. In this regime, both single-particle minima–corresponding to the dressed spin states mentioned in the introduction–are equally populated, resulting in a configuration with zero spin polarization $\langle\sigma_{z}\rangle=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\sigma_{z}\Psi$. This balanced occupation leads to the formation of spatial modulations in the density profile. The emergence of such modulations indicates spontaneous breaking of translation symmetry. When combined with the global $\mathrm{U}(1)$ phase symmetry breaking intrinsic to Bose-Einstein condensation, this reveals the supersolid nature of the stripe phase. This interpretation is further corroborated by the properties of the Bogoliubov spectrum [76, 58] (see also Sec. V).
In addition to the stripe phase, which is the main focus of this paper, the equilibrium phase diagram of spin-orbit-coupled BECs includes two non-supersolid phases [69, 70]. The plane-wave phase corresponds to a scenario where atoms occupy only one of the two single-particle energy minima. Depending on which minimum is chosen, the momentum and spin polarization can take opposite values: $\langle\bm{p}\rangle=\pm N\hbar k_{1}^{\mathrm{PW}}\hat{\bm{e}}_{x}$, $\langle\sigma_{z}\rangle=\pm Nk_{1}^{\mathrm{PW}}/k_{R}$, where
$k_{1}^{\mathrm{PW}}=k_{R}\sqrt{1-\left(\frac{\Omega_{R}}{\Omega_{\mathrm{cr2}}}\right)^{2}}$ (6)
and $\Omega_{\mathrm{cr2}}$ is the critical Raman coupling defined below. In the plane-wave phase, the translation symmetry is unbroken, and the density remains uniform and equal to the average density $\bar{n}=N/V$. This feature is shared by the second non-supersolid phase, the single-minimum phase in which both the momentum and spin polarization are zero.
The non-supersolid phases appear at higher Raman coupling values, where the energy cost of density modulations becomes prohibitively high. At low average density $\bar{n}$, the stripe and plane-wave phases are separated by a first-order transition, which occurs at a critical Raman coupling $\hbar\Omega_{\mathrm{cr1}}=4E_{R}\sqrt{2g_{ss}/(g_{dd}+2g_{ss})}$ in the $\bar{n}\to 0$ limit [69, 70]. Conversely, the second-order transition from the plane-wave to the single-minimum phase takes place at the larger critical value $\hbar\Omega_{\mathrm{cr2}}=2(2E_{R}-g_{ss}\bar{n})$. As $\bar{n}$ increases, the plane-wave phase becomes less favorable and eventually disappears. As a result, the system can have a first order transition directly from the stripe phase to the single-minimum phase [70, 77].
III Traveling wave patterns
We will now demonstrate how to construct solutions to the Gross-Pitaevskii equation (1) that describe stripe patterns moving with constant velocity. This process will be illustrated in both the laboratory (Sec. III.1) and comoving (Sec. III.2) frames. The laboratory frame corresponds to the rest frame of the Raman lasers responsible for generating the spin-orbit coupling, while the comoving frame is defined as the rest frame of the crystal pattern. We will show that, despite the lack of Galilean invariance in spin-orbit-coupled BECs, these two frames provide equivalent descriptions. Finally, in Sec. III.3, we will explore the key properties of the order parameter associated with the moving stripes.
III.1 Laboratory frame
Consider a periodic wave propagating in a spin-orbit-coupled BEC with a constant velocity $\bm{v}$ in the laboratory frame. The order parameter takes the standard form
$\Psi(\bm{r},t)=\mathrm{e}^{-\mathrm{i}\mu t/\hbar}\Psi_{0}(\bm{r}-\bm{v}t)\,,$ (7)
where $\mu$ is the chemical potential in the laboratory frame, and $\Psi_{0}$ is a function satisfying periodic boundary conditions in all three spatial directions. Substituting the traveling-wave Ansatz (7) into the Gross-Pitaevskii equation (1) and transforming to the comoving frame coordinates $\bm{r}^{\prime}=\bm{r}-\bm{v}t$ and $t^{\prime}=t$ (note that throughout this work, primed quantities refer to the comoving frame), we obtain the following time-independent equation for $\Psi_{0}$:
$\begin{split}&{}h_{\mathrm{SO}}\Psi_{0}+g_{dd}\left(\Psi_{0}^{\dagger}\Psi_{0}\right)\Psi_{0}+g_{ss}\left(\Psi_{0}^{\dagger}\sigma_{z}\Psi_{0}\right)\sigma_{z}\Psi_{0}\\ &{}=(\mu+\bm{v}\cdot\bm{p})\Psi_{0}\,,\end{split}$ (8)
where now $\bm{p}=-\mathrm{i}\hbar\nabla_{\bm{r}^{\prime}}$. For $\bm{v}=0$, Eq. (8) reduces to the standard time-independent Gross-Pitaevskii equation, whose solutions represent the stationary states of the condensate, including the ground state. The $\bm{v}$-dependent term on the right hand side was previously included in Ref. [61] to construct supercurrent states, though that study considered only uniform plane-wave configurations. Note that Eq. (8) corresponds to the stationarity condition of the functional
$E_{\mathrm{GC}}^{\prime}=E-\mu N-\bm{v}\cdot\langle\bm{p}\rangle\,,$ (9)
which is the grand canonical energy in the comoving frame (as detailed below). Therefore, $\mu$ and $\bm{v}$ can be interpreted as Lagrange multipliers, determined by fixing the values of the particle number $N$ and the momentum $\langle\bm{p}\rangle$.
III.2 Comoving frame
The order parameter $\Psi^{\prime}(\bm{r}^{\prime},t^{\prime})$ in the comoving frame satisfies the time-dependent Gross-Pitaevskii equation
$\mathrm{i}\hbar\partial_{t^{\prime}}\Psi^{\prime}=h_{\mathrm{SO}}^{\prime}\Psi^{\prime}+g_{dd}\left(\Psi^{\prime\dagger}\Psi^{\prime}\right)\Psi^{\prime}+g_{ss}\left(\Psi^{\prime\dagger}\sigma_{z}\Psi^{\prime}\right)\sigma_{z}\Psi^{\prime}\,.$ (10)
This equation differs from its counterpart in the laboratory frame, Eq. (8), due to an additional detuning term arising from the transformed spin-orbit Hamiltonian,
$h_{\mathrm{SO}}^{\prime}=h_{\mathrm{SO}}-v_{x}\hbar k_{R}\sigma_{z}\,.$ (11)
This extra term can be understood by noting that in the comoving frame the Raman lasers move with velocity $-\bm{v}$. As a result, the frequency of the light field experienced by the atoms undergoes a Doppler shift of $-2v_{x}k_{R}$, which leads to the corresponding shift in the Raman detuning $\delta_{R}$ by the same amount [60, 78].
In the comoving frame, the periodic wave pattern is stationary, while the superfluid background is boosted with velocity $-\bm{v}$. This configuration can be described by a stationary wave function of the form
$\Psi^{\prime}(\bm{r}^{\prime},t^{\prime})=\mathrm{e}^{-\mathrm{i}\mu^{\prime}t^{\prime}/\hbar}\mathrm{e}^{-\mathrm{i}m\bm{v}\cdot\bm{r}^{\prime}/\hbar}\Psi_{0}(\bm{r}^{\prime})\,.$ (12)
Substituting this Ansatz into Eq. (10) and setting $\mu^{\prime}=\mu+mv^{2}/2$, one can easily recover Eq. (8). This confirms that the function $\Psi_{0}$ introduced in Eq. (12) is identical to the one used in the traveling-wave Ansatz (7). Additionally, the relation between the energies in the two frames is given by
$E^{\prime}=E-\bm{v}\cdot\langle\bm{p}\rangle+\frac{Nmv^{2}}{2}\,.$ (13)
Taking the derivative of both sides of this equation with respect to $N$, we see that $\mu^{\prime}$ is indeed the chemical potential in the comoving frame. Furthermore, the grand canonical energy $E_{\mathrm{GC}}^{\prime}=E^{\prime}-\mu^{\prime}N$ reproduces Eq. (9) above.
III.3 Traveling stripe Ansatz
The full condensate order parameter in the supersolid stripe phase can be expressed as an expansion of the form [76, 58]
$\Psi_{0}(\bm{r}^{\prime})=\sqrt{\bar{n}}\sum_{\text{{\hbox{\bar{m}}} odd}}\tilde{\Psi}_{\bar{m}}\mathrm{e}^{\mathrm{i}\bar{m}k_{1}x^{\prime}}\,.$ (14)
In this expression, the $\tilde{\Psi}_{\bar{m}}$’s are two-component expansion coefficients, and the sum runs over both positive and negative odd integers. The presence of the $\bar{m}=\pm 1$ terms in the Ansatz indicates the simultaneous occupation of the two minima of the single-particle spectrum. In contrast, the $|\bar{m}|>1$ harmonics arise due to the nonlinearity of the Gross-Pitaevskii equation (8) that $\Psi_{0}$ satisfies. Note that the order parameter (14) is $\pi/k_{1}$-antiperiodic (and therefore $2\pi/k_{1}$-periodic) in $x^{\prime}$, while it is constant in $y^{\prime}$ and $z^{\prime}$ due to the absence of spin-orbit coupling along these two directions. The quantity $k_{1}$ represents the half-length of the stripe wave vector, $2k_{1}\hat{\bm{e}}_{x}$. Its optimal value, along with the expansion coefficients $\tilde{\Psi}_{\bar{m}}$, is determined by minimizing the energy (3) for an order parameter of the form (14). This leads to the following expression for $k_{1}$ as a function of the $\tilde{\Psi}_{\bar{m}}$’s [58]:
$k_{1}=k_{R}\frac{\sum_{\text{{\hbox{\bar{m}}} odd}}\bar{m}\tilde{\Psi}_{\bar{m}}^{\dagger}\sigma_{z}\tilde{\Psi}_{\bar{m}}}{\sum_{\text{{\hbox{\bar{m}}} odd}}\bar{m}^{2}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}}\,.$ (15)
Typically, the optimal $k_{1}$ is found to be approximately equal to the value $k_{1}^{\mathrm{SP}}$ obtained within the single-particle model, with corrections due to interparticle interactions [70, 58]. One can also explore configurations with non-optimal $k_{1}$, whose stability has been analyzed in both the supersolid [79] and non-supersolid [61] regimes.
In the previous work on the stationary stripe phase [76, 58] the energy minimization procedure was carried out under the constraint of fixed particle number (4), which leads to the normalization condition
¯$\sum_{\text{{\hbox{\bar{m}}} odd}}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}=1$ (16)
for the expansion coefficients. Regarding the momentum (5), the independence of the order parameter (14) from $y$ and $z$ immediately implies that the corresponding components $\langle p_{y}\rangle$ and $\langle p_{z}\rangle$ vanish. On the other hand, the result $\langle p_{x}\rangle=0$ in the stationary stripe phase can be understood through more subtle symmetry considerations. The time-independent Gross-Pitaevskii equation (8) with $\bm{v}=0$ is invariant under the action of the two operators $\sigma_{x}\mathcal{P}$ and $\sigma_{z}\mathcal{T}$, which combine spin rotation with parity ($\mathcal{P}$) and time reversal ($\mathcal{T}$), respectively. [^1] In a spin-$1/2$ system, the time-reversal operator is $\mathcal{T}=\mathrm{i}\sigma_{y}\mathcal{K}$ where $\mathcal{K}$ denotes complex conjugation. Thus, the order parameter resulting from the application of either of these two operators to Eq. (14) remains a solution of Eq. (8) and can differ from the original only by a global phase rotation and spatial translation. In terms of the expansion coefficients, this means that $\tilde{\Psi}_{-\bar{m}}^{*}=\sigma_{x}\tilde{\Psi}_{\bar{m}}$ up to a phase factor [76, 58]. Using this property, one can easily show that $\langle p_{x}\rangle$ vanishes when $\bm{v}=0$. Conversely, when the stripes are in motion, the action of $\sigma_{x}\mathcal{P}$ and $\sigma_{z}\mathcal{T}$ on Eq. (14) generates a stripe pattern moving with velocity $-\bm{v}$. In this case, the $x$ component of the momentum (5) is nonzero and is related to the expansion coefficients $\tilde{\Psi}_{\bar{m}}$ as follows:
¯$\sum_{\text{{\hbox{\bar{m}}} odd}}\bar{m}\tilde{\Psi}_{\bar{m}}^{\dagger}\tilde{\Psi}_{\bar{m}}=\frac{\langle p_{x}\rangle}{N\hbar k_{1}}\equiv s_{p}\,.$ (17)
Neglecting the typically small contribution from the $|\bar{m}|>1$ harmonics in Eq. (14), one can interpret $s_{p}$ as the imbalance between the occupations of the two single-particle energy minima. For a Bloch-wave-like order parameter of the form (14), the momentum per particle is constrained to the first Brillouin zone, meaning that $-\hbar k_{1}\leq\langle p_{x}\rangle/N\leq\hbar k_{1}$, which leads to the condition $-1\leq s_{p}\leq 1$. As we will see, $s_{p}$ in our spin-orbit-coupled BEC plays a role similar to that of the spin imbalance $\langle\sigma_{z}\rangle/N$ in the absence of spin-orbit coupling. In this analogy, $s_{p}=0$ corresponds to a balanced coherent mixture of the two degenerate minima of the single-particle spectrum, which defines the stripe phase at equilibrium. On the other hand, we will show that $s_{p}=1$ ($s_{p}=-1$) corresponds to a fully polarized state in momentum space, where all the atoms condense into the single-particle ground state with positive (negative) momentum. This is nothing else than the well-known plane-wave phase (see Sec. II). Interestingly, $s_{p}$ and $\langle\sigma_{z}\rangle/N$ coincide in the limit of zero Raman coupling, where the single-particle energy minima are characterized by well-defined spin (see Sec. IV.1). The explicit determination of the order parameter (14) for moving stripe patterns and the related observables will be discussed in the next section.
IV Properties of moving striped patterns
After establishing the general framework, we now proceed with a detailed analysis of the moving stripe patterns. We begin by examining the exact solution for the case of zero Raman coupling, which can be mapped to an unbalanced BEC mixture (Sec. IV.1). Next, in Sec. IV.2, we describe the numerical and perturbative methods used to calculate the order parameter at finite Raman coupling. Finally, the results for several key observables are presented in Sec. IV.3.
IV.1 Solution at zero Raman coupling. Unbalanced mixture
At zero Raman coupling ($\Omega_{R}=0$), the lowest-energy solution of Eq. (8) that satisfies the constraints (16) and (17) for a given $s_{p}$ has a simple analytical form:
$\Psi_{0}^{(0)}(\bm{r}^{\prime})=\sqrt{\bar{n}}\left[\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(0)}\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}k_{R}x^{\prime}}+\begin{pmatrix}0\\ \tilde{\Psi}_{-1,\downarrow}^{(0)}\end{pmatrix}\mathrm{e}^{-\mathrm{i}k_{R}x^{\prime}}\right]\,,$ (18)
where $\tilde{\Psi}_{+1,\uparrow}^{(0)}=\sqrt{(1+s_{p})/2}\,\mathrm{e}^{\mathrm{i}(\theta+\Delta\theta/2)}$ and $\tilde{\Psi}_{-1,\downarrow}^{(0)}=\sqrt{(1-s_{p})/2}\,\mathrm{e}^{\mathrm{i}(\theta-\Delta\theta/2)}$. This solution has $\mu=G_{dd}$ and $\Delta\mu/2\equiv\hbar k_{R}v_{x}=G_{ss}s_{p}$, where $G_{dd}=g_{dd}\bar{n}$ and $G_{ss}=g_{ss}\bar{n}$. Equation (18) represents a special case of the stripe Ansatz (14), where only the $\bar{m}=\pm 1$ terms are nonzero, and the wave vector takes the value $k_{1}=k_{R}$, which results from energy minimization. In contrast, the global phase of the order parameter, $\theta$, and the relative phase between the two spatially oscillating terms, $\Delta\theta$, are not determined by energy minimization and can be chosen arbitrarily. The energy per particle for the configuration (18) in the laboratory frame is $E/N=(G_{dd}+G_{ss}s_{p}^{2})/2$. The relation $\langle\bm{p}\rangle=\hbar k_{R}\langle\sigma_{z}\rangle\hat{\bm{e}}_{x}$ holding at zero $\Omega_{R}$ implies that in this case, $s_{p}$ can be used to quantify both the spin polarization, $\langle\sigma_{z}\rangle=Ns_{p}$, and the momentum along $x$, $\langle p_{x}\rangle=N\hbar k_{R}s_{p}$. We emphasize that only the latter physical interpretation remains valid at finite Raman coupling [see Eq. (17)].
Equation (18) describes a standard two-component BEC mixture with uniform density $\bar{n}$ and spin imbalance $\langle\sigma_{z}\rangle/N=s_{p}$. This becomes even more evident when considering the full order parameter in the laboratory frame, given by $\mathrm{e}^{-\mathrm{i}\mu t/\hbar}\Psi_{0}^{(0)}(\bm{r}-\bm{v}t)$ [see Eq. (7)], and performing the space-dependent spin rotation $\exp(-\mathrm{i}k_{R}\sigma_{z}x)$. This procedure transforms the order parameter into the canonical space-independent form [^2] Applying the unitary transformation $\exp(-\mathrm{i}k_{R}\sigma_{z}x)$ to the single-particle Hamiltonian (2) with $\Omega_{R}=\delta_{R}=0$ reduces it to the standard kinetic energy $\bm{p}^{2}/2m$. Consequently, Eq. (1) becomes the Gross-Pitaevskii equation for a two-component BEC without spin-orbit coupling, and its uniform solutions take the form in Eq. (19).
$\Psi_{\mathrm{mix}}(t)=\sqrt{\bar{n}}\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(0)}\mathrm{e}^{-\mathrm{i}\mu_{\uparrow}t/\hbar}\\ \tilde{\Psi}_{-1,\downarrow}^{(0)}\mathrm{e}^{-\mathrm{i}\mu_{\downarrow}t/\hbar}\end{pmatrix}\,.$ (19)
Notice that the combinations $\mu_{\uparrow,\downarrow}=\mu\pm\Delta\mu/2=G_{dd}\pm G_{ss}s_{p}$ correspond exactly to the chemical potentials of the two components in a binary BEC mixture with equal masses and symmetric intraspecies interactions [80, 81]. The global phase of the order parameter (19), $\theta-\mu t/\hbar$, evolves in time at a rate determined by the average chemical potential, $\mu=(\mu_{\uparrow}+\mu_{\downarrow})/2$. The freedom in choosing the value of $\theta$ at $t=0$ reflects the spontaneous breaking of global $\mathrm{U}(1)$ phase symmetry, a hallmark of Bose-Einstein condensation. The polarization vector of the mixture (19) is defined as $\langle\bm{\sigma}\rangle_{\mathrm{mix}}=\left(\langle\sigma_{x}\rangle_{\mathrm{mix}},\langle\sigma_{y}\rangle_{\mathrm{mix}},\langle\sigma_{z}\rangle_{\mathrm{mix}}\right)$, where
$\langle\sigma_{x}\rangle_{\mathrm{mix}}$ ${}=N\sqrt{1-s_{p}^{2}}\cos(\Delta\mu\,t/\hbar-\Delta\theta)\,,$ (20a)
$\langle\sigma_{y}\rangle_{\mathrm{mix}}$ ${}=N\sqrt{1-s_{p}^{2}}\sin(\Delta\mu\,t/\hbar-\Delta\theta)\,,$ (20b)
$\langle\sigma_{z}\rangle_{\mathrm{mix}}$ ${}=Ns_{p}\,,$ (20c)
and the subscript “mix” indicates that expectation values are evaluated with the order parameter (19). When $s_{p}\neq\pm 1$, the polarization vector exhibits a nonzero component in the $xy$ plane, signaling the spontaneous breaking of spin rotational symmetry around the $z$ axis. The azimuthal angle of the polarization is equal to minus the relative phase $\Delta\theta-\Delta\mu\,t/\hbar$ between the two components of the mixture (19). At $t=0$, the direction of the in-plane polarization is set by the arbitrary parameter $-\Delta\theta$, and it subsequently rotates with a frequency determined by the chemical potential difference $\Delta\mu=\mu_{\uparrow}-\mu_{\downarrow}$. This results in a precession of $\langle\bm{\sigma}\rangle_{\mathrm{mix}}$ around the $z$ axis, a phenomenon known in polariton physics as self-induced Larmor precession [82]. This behavior originates from the $g_{ss}$-dependent term in the Gross-Pitaevskii equation (1), which effectively acts as a self-induced Zeeman splitting.
IV.2 Evaluation of the moving stripe order parameter
The discussion in the previous section demonstrates that, although the order parameter at zero Raman coupling $\Omega_{R}$, given by Eq. (18), formally breaks translation invariance, this has no physical implications, since the spatial dependence can be eliminated via a unitary transformation. However, the situation changes dramatically once the Raman coupling is introduced. On the one hand, due to the $\sigma_{x}$-dependent term, the breaking of spin rotational symmetry around the $z$ axis becomes explicit. On the other hand, the spontaneous breaking of translation symmetry now has observable consequences, most notably the emergence of density fringes. In this context, the self-induced Larmor precession characteristic of the zero spin-orbit coupling regime is replaced at finite $\Omega_{R}$ by the self-induced translational motion of the stripe pattern.
In the present work, we employ two complementary approaches to determine the order parameter of traveling stripe patterns. The first is a numerical method, which begins by inserting the Ansatz (14) into the total energy (3) and performing the spatial integration. This yields an expression for the energy as a function of the wave vector $k_{1}$ and the expansion coefficients $\tilde{\Psi}_{\bar{m}}$. The energy is then minimized under two constraints enforced via Lagrange multipliers: the chemical potential $\mu$ associated with the normalization condition (16), and the quantity $\Delta\mu/2\equiv\hbar k_{1}v_{x}$ which fixes the value of the momentum (more precisely, the dimensionless ratio $s_{p}=\langle p_{x}\rangle/N\hbar k_{1}$) as required by Eq. (17). Analogously to the zero-$\Omega_{R}$ case discussed in Sec. IV.1, $\Delta\mu$ can here be interpreted as a chemical potential difference between the two minima of the single-particle spectrum and is directly related to the fringe velocity $v_{x}$, while $\mu$ represents the average chemical potential. The stationarity condition of the energy per particle with respect to $k_{1}$, taken at fixed $\mu$, $\Delta\mu$, and $\tilde{\Psi}_{\bar{m}}$, yields the same expression for the optimal wave vector as in the $\bm{v}=0$ case, namely, Eq. (15[58]. To perform the numerical minimization over the components of the $\tilde{\Psi}_{\bar{m}}$’s, the expansion (14) must be truncated to a finite number of harmonics. In our calculations, we retain terms with $-9\leq\bar{m}\leq 9$, which is sufficient to achieve high accuracy in all physical quantities of interest. This choice is validated by direct comparison with the numerical solution of the Gross-Pitaevskii equation (8) in a periodic box, showing excellent agreement.
In Ref. [58], a perturbative approach was developed to analytically compute the equilibrium order parameter in the stripe phase, along with the associated observables. This method has proven highly accurate deeply in the double-minimum regime of the single-particle spectrum, which corresponds to low Raman coupling, i.e., $\hbar\Omega_{R}/4E_{R}\ll 1$. In the present work, we extend this perturbative approach to investigate the case of a moving striped condensate, thereby gaining deeper insight into the underlying physics. This method serves as an analytical alternative to the numerical minimization discussed earlier. Notably, the two approaches yield excellent agreement for small enough values of the Raman coupling (see next section for comparison). We begin by considering the power series expansion of the order parameter,
$\Psi_{0}(\bm{r}^{\prime})=\Psi_{0}^{(0)}(\bm{r}^{\prime})+\sum_{l=1}^{+\infty}\Psi_{0}^{(l)}(\bm{r}^{\prime})\,,$ (21)
along with the expansions for the average chemical potential,
$\mu=G_{dd}+\sum_{l=1}^{+\infty}\mu^{(l)}\,,$ (22)
the chemical potential semi-difference,
$\frac{\Delta\mu}{2}=G_{ss}s_{p}+\sum_{l=1}^{+\infty}\frac{\Delta\mu^{(l)}}{2}\,,$ (23)
and the wave vector,
$k_{1}=k_{R}+\sum_{l=1}^{+\infty}k_{1}^{(l)}\,.$ (24)
Here and in the following, the superscript “$(l)$” denotes the contribution of order $l$ in the small expansion parameter $\hbar\Omega_{R}/4E_{R}$. The zero order terms in each series coincide with the results obtained in the absence of Raman coupling, as discussed in Sec. IV.1. Substituting these expansions into the stationary Gross-Pitaevskii equation (8) yields a set of recurrence relations, which allow for the computation of each order-$l$ correction in terms of lower-order results up to $l-1$. A detailed derivation of this procedure, together with the expressions for the order parameter up to the second order in $\hbar\Omega_{R}/4E_{R}$, is given in Appendix A and Appendix B.
Figure 1: Density profiles of spin-orbit-coupled BEC in the stripe phase for (a) $s_{p}=0$ (stationary stripes), (b) $s_{p}=0.5$, and (c) $s_{p}=0.99$. The numerically obtained profiles of the spin-up (blue solid lines) and spin-down (yellow dashed lines) components are shown. Close to each curve the prediction of the perturbative approach is also plotted (black dotted lines). For definiteness, the fringe offset has been chosen such that a density minimum occurs at $x=0$. The other parameters are $\hbar\Omega_{R}/E_{R}=1.0$, $G_{dd}/E_{R}=1.2$, and $G_{ss}/E_{R}=0.32$.
IV.3 Results for the observables
After evaluating the condensate order parameter using the two methods outlined in the previous section, we now discuss the behavior of several key observables. Throughout this and the following sections, we set the interaction parameters to $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$ consistent with the values used in previous works [76, 58]. With this choice of parameters, the first-order transition separating the stripe and plane-wave phases in the ground-state phase diagram occurs at the critical Raman coupling $\hbar\Omega_{\mathrm{cr1}}/E_{R}=2.70$. However, the stripe phase remains metastable up to the spinodal point $\hbar\Omega_{\mathrm{sp}}/E_{R}=2.85$ [58]. In most figures of this work, we compare results for two different values of the Raman coupling. The smaller value, $\hbar\Omega_{R}/E_{R}=1.0$, lies within the range of validity for the perturbative approach, where the predictions of this method match the numerical results excellently. The other value, $\hbar\Omega_{R}/E_{R}=2.6$, is so large that only the numerical method is applicable. It is important to note that the interaction parameters used here differ significantly from those in the original experiment of Ref. [68]. The states of 87Rb employed in that experiment are indeed characterized by a very low degree of miscibility, as evidenced by the small ratio $g_{ss}/g_{dd}\sim 10^{-3}$, resulting in a small critical Raman coupling $\hbar\Omega_{\mathrm{cr1}}/E_{R}=0.19$. This limits the maximum achievable contrast of the stripes and thus restricts the possibility of observing significant supersolidity effects. To address this limitation, potential strategies include using atomic species with tunable interactions, as demonstrated in [22] employing 41K atoms. Alternatively, one could reduce the spatial overlap between the two spin components along a direction parallel to the stripes. This could be achieved by applying a quasi-two-dimensional spin-dependent trapping potential [83, 84, 85] or by implementing spin-orbit coupling between two orbital states within a superlattice potential [86].
We first investigate the effect of stripe motion on the profiles of the total density and spin density. The total density is expressed as
$n(\bm{r}^{\prime})=\bar{n}+\sum_{\bar{m}=1}^{+\infty}\tilde{n}_{\bar{m}}\cos[\bar{m}(2k_{1}x^{\prime}+\Delta\theta)]$ (25)
and the spin density is given by
$s_{z}(\bm{r}^{\prime})=\frac{\langle\sigma_{z}\rangle}{V}+\sum_{\bar{m}=1}^{+\infty}\tilde{s}_{z,\bar{m}}\cos[\bar{m}(2k_{1}x^{\prime}+\Delta\theta)]\,.$ (26)
As with the order parameter in Eq. (14), the total and spin densities are represented as Fourier series. These densities oscillate around their respective average values, $\bar{n}$ and $\langle\sigma_{z}\rangle/V$. The phase $\Delta\theta$ sets the offset of the density fringes. Like in the case of $\Omega_{R}=0$ (see Sec. IV.1), $\Delta\theta$ is randomly selected by the system following the spontaneous breaking of translation invariance. The expansion coefficients $\tilde{n}_{\bar{m}}$ and $\tilde{s}_{z,\bar{m}}$ can be derived directly from those of the order parameter (and the same applies for the spin polarization $\langle\sigma_{z}\rangle$, which is discussed below). At second order in the Raman coupling, only the terms with $\bar{m}=1$ and $\bar{m}=2$ are nonzero. These coefficients are given by
$\frac{\tilde{n}_{1}}{\bar{n}}={}$ ${}-\frac{8E_{R}^{3}(2E_{R}+G_{ss})\sqrt{1-s_{p}^{2}}}{D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,,$ (27a)
$\frac{\tilde{n}_{2}}{\bar{n}}={}$ ${}-\frac{2E_{R}^{2}(1-s_{p}^{2})N_{\tilde{n}_{2}}}{D_{1}^{2}D_{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,$ (27b)
$\begin{split}\frac{\tilde{s}_{z,1}}{\bar{n}}={}&{}2E_{R}\big{[}4E_{R}^{2}G_{ss}+2E_{R}\left(2E_{R}+G_{ss}\right)G_{dd}\\ {}&{}-G_{ss}^{3}s_{p}^{2}\big{]}\frac{s_{p}\sqrt{1-s_{p}^{2}}}{D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,,\end{split}$ (27c)
$\frac{\tilde{s}_{z,2}}{\bar{n}}={}$ ${}\frac{E_{R}^{2}s_{p}(1-s_{p}^{2})N_{\tilde{s}_{z,2}}}{D_{1}^{2}D_{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,.$ (27d)
Here and in the other perturbative formulas presented in this section, the quantities $D_{1}$ and $D_{2}$ appearing in the denominators are even polynomial functions of $s_{p}$. Their full expressions are provided in Appendix B, see Eqs. (59) and (60). The same applies to the numerators $N_{\tilde{n}_{2}}$ and $N_{\tilde{s}_{z,2}}$, given in Eqs. (61) and (62), respectively. It is worth noting that the total density $n$ is an even function of $s_{p}$, while the spin density $s_{z}$ is an odd function. In Fig. 1, we show the combinations $n_{\uparrow,\downarrow}=(n\pm s_{z})/2$, corresponding to the densities of the two spin components, for various values of $s_{p}$. When $s_{p}=0$ [Fig. 1(a)], the spin density vanishes, yielding equal profiles for the two components: $n_{\uparrow}=n_{\downarrow}$. For $s_{p}=0.5$ [Fig. 1(b)], the system becomes both globally and locally spin-polarized, with the two components oscillating around different average values, but remaining in phase. Notably, the minority component (spin-down for $s_{p}>0$, spin-up for $s_{p}<0$) exhibits a larger oscillation amplitude than the majority one. Finally, for $s_{p}=0.99$ [Fig. 1(c)], both components display only weak density modulations around their mean values which closely resemble those of the uniform plane-wave phase.
Figure 2: The fringe contrast [(a1)-(a2)], the fringe wave vector [(b1)-(b2)], and the spin polarization [(c1)-(c2)] as functions of $s_{p}$. In panels [(a1)-(a2)], we show the fringe contrast of the total density (red), as well as the contrast of the majority (blue) and minority (yellow) spin components. Each observable is plotted for two different values of the Raman coupling: $\hbar\Omega_{R}/E_{R}=1.0$ (left panels) and $\hbar\Omega_{R}/E_{R}=2.6$ (right panels). Solid lines correspond to the numerical results. In the left column, close to each of these solid lines, we add a black dotted line showing the predictions of the perturbative approach. The interaction parameters are the same as in Fig. 1: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
To gain a deeper understanding of the density profiles, it is useful to explicitly evaluate their contrast. For the total density $n$, the contrast is defined as
$\mathcal{C}=\frac{n_{\mathrm{max}}-n_{\mathrm{min}}}{n_{\mathrm{max}}+n_{\mathrm{min}}}\,,$ (28)
where $n_{\mathrm{max}}$ and $n_{\mathrm{min}}$ denote the maximum and minimum values of $n$ within one oscillation period, respectively. Generally, the contrast is an odd function of the Raman coupling $\Omega_{R}$ since changing the sign of $\Omega_{R}$ swaps the positions of the density maxima and minima. Expressing the contrast in terms of the coefficients of the Fourier expansion (25) gives
$\mathcal{C}=\frac{\left|\sum_{\bar{m}=1}^{+\infty}\tilde{n}_{2\bar{m}-1}\right|}{\bar{n}+\sum_{\bar{m}=1}^{+\infty}\tilde{n}_{2\bar{m}}}\,.$ (29)
At leading order, the contrast depends linearly on $\Omega_{R}$ and matches (up to a sign) the expression given in Eq. (27a). Figures 2(a1)-(a2) show $\mathcal{C}$ as a function of $s_{p}$ for fixed $\Omega_{R}$ and interaction parameters. The contrast reaches its maximum at $s_{p}=0$, where it coincides with the value reported in Ref. [58]. As $s_{p}$ increases, the contrast decreases, indicating that the density modulations weaken as the stripes move. The contrast vanishes at $s_{p}=\pm 1$, where the fringes disappear entirely and the system becomes a uniform plane-wave condensate. Comparison of the two panels reveals that, similarly to the stationary case, the contrast grows with increasing Raman coupling even at finite $s_{p}$.
The contrast of the density profiles of the individual spin components is defined analogously to Eq. (28). As shown in Figs. 2(a1)-(a2), the contrast $\mathcal{C}_{>}$ of the majority component–like that of the total density–is maximal at $s_{p}=0$, decreases monotonically with increasing $|s_{p}|$, and vanishes at $|s_{p}|=1$. In contrast, the behavior of the minority component contrast $\mathcal{C}_{<}$ is nonmonotonic. It initially increases with $|s_{p}|$, reaching the maximum value of $1$ (sooner for larger Raman coupling), and then rapidly drops to zero as $|s_{p}|\to 1$. At first order in the Raman coupling, $\mathcal{C}_{>}$ and $\mathcal{C}_{<}$ can be directly expressed in terms of the coefficients of the expansions (25) and (26). For $s_{p}>0$, where the majority (minority) component corresponds to spin-up (spin-down), one finds $\mathcal{C}_{>}=-(\tilde{n}_{1}+\tilde{s}_{z,1})/[\bar{n}(1+s_{p})]$ and $\mathcal{C}_{<}=-(\tilde{n}_{1}-\tilde{s}_{z,1})/[\bar{n}(1-s_{p})]$. For $s_{p}<0$, the roles of the two components are interchanged. From Figs. 2(a1)-(a2) we observe that the perturbative prediction for $\mathcal{C}_{>}$ matches the numerical results very accurately over the whole range of $|s_{p}|$. However for $\mathcal{C}_{<}$ the perturbative expression is reliable only at small $|s_{p}|$. It eventually diverges unphysically as $|s_{p}|\to 1$. This divergence arises because in the limit $|s_{p}|\to 1$ and for vanishing Raman coupling the denominator in the definition of the contrast [Eq. (28)] tends to zero for the minority component. Consequently, the perturbative approach becomes invalid for calculating $\mathcal{C}_{<}$ at large $|s_{p}|$.
The wave vector $k_{1}$, which determines the periodicity of the stripe phase, also exhibits a dependence on $s_{p}$, albeit much more gradually than the contrast does [see Figs. 2(b1)-(b2)]. This quantity is primarily governed by single-particle physics, with only minor corrections arising from interactions [70, 58]. By substituting the expressions derived in Appendixes into Eq. (15) one obtains the perturbative formula
$k_{1}=k_{R}\left[1-\frac{2E_{R}^{2}N_{k_{1}}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,,$ (30)
where the numerator $N_{k_{1}}$ is given in Eq. (63). One can see that $k_{1}$ smoothly interpolates between its value in the stationary stripe phase at $s_{p}=0$ [58] and its value in the plane-wave phase at $|s_{p}|=1$, denoted by $k_{1}^{\mathrm{PW}}$, Eq. (6). Notably, the perturbative result (30) when evaluated at $|s_{p}|=1$ exactly matches the second order expansion of $k_{1}^{\mathrm{PW}}$ in $\hbar\Omega_{R}/4E_{R}$. Finally, we emphasize that $k_{1}$, like the other observables discussed below, should remain invariant under the sign change of the Raman coupling. It is therefore an even function of $\Omega_{R}$.
As already illustrated in Fig. 1, moving stripes exhibit a population imbalance between the spin-up and spin-down components, resulting in a finite spin polarization. In general, the spin polarization $\langle\sigma_{z}\rangle$ is an odd function of $s_{p}$, being positive for $s_{p}>0$ and negative for $s_{p}<0$. Within second order perturbation theory in the Raman coupling, it takes the form
$\langle\sigma_{z}\rangle=Ns_{p}\left[1-\frac{2E_{R}^{2}(2E_{R}+G_{ss})N_{\langle\sigma_{z}\rangle}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,,$ (31)
where the coefficient $N_{\langle\sigma_{z}\rangle}$ is defined in Eq. (64). This expression shows that $\langle\sigma_{z}\rangle$ increases linearly with $s_{p}$ for small $|s_{p}|$, in agreement with numerical results plotted in Figs. 2(c1)-(c2). The spin polarization reaches its maximum magnitude at $s_{p}=\pm 1$, where it coincides with the value $Nk_{1}^{\mathrm{PW}}/k_{R}$ characteristic of the plane-wave phase (see Sec. II). It is worth noting that $\langle\sigma_{z}\rangle$, like the stripe wave vector $k_{1}$, decreases as the Raman coupling increases. This behavior reflects the reduced separation between the single-particle energy minima at stronger Raman couplings, which in turn leads to a diminished spin polarization associated with each minimum.
Figure 3: The energy per particle [(a1)-(a2)], the chemical potential [(b1)-(b2)], the chemical potential difference [(c1)-(c2)], and the fringe velocity [(d1)-(d2)] as functions of $s_{p}$. As in Fig. 2, each quantity is shown for two different values of the Raman coupling: $\hbar\Omega_{R}/E_{R}=1.0$ (left panels) and $\hbar\Omega_{R}/E_{R}=2.6$ (right panels). Solid lines represent the results of numerical calculations. In the left panels, close to each of these lines, we add a black dotted line indicating the perturbative predictions. The interaction parameters are the same as in the previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
The perturbative approach also allows for the derivation of analytical expressions for key thermodynamic quantities. The energy per particle, obtained from Eq. (3), reads
$\frac{E}{N}=\frac{G_{dd}}{2}+\frac{G_{ss}}{2}s_{p}^{2}-\frac{E_{R}^{2}N_{E}}{D_{1}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,$ (32)
while the average chemical potential is given by
$\mu=G_{dd}+\frac{2E_{R}^{2}N_{\mu}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,$ (33)
and the chemical potential semi-difference takes the form
$\frac{\Delta\mu}{2}=s_{p}\left[G_{ss}-\frac{2E_{R}^{2}(2E_{R}+G_{ss})N_{\Delta\mu}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,.$ (34)
The coefficients $N_{E}$, $N_{\mu}$, and $N_{\Delta\mu}$ are given in Eqs. (65), (66), and (67), respectively. These expressions satisfy the thermodynamic relations
$\mu=\left(\frac{\partial E}{\partial N}\right)_{S_{p}}\,,\quad\frac{\Delta\mu}{2}=\left(\frac{\partial E}{\partial S_{p}}\right)_{N}\,,$ (35)
where $S_{p}=Ns_{p}$. Using the identity $v_{x}=\Delta\mu/2\hbar k_{1}$, and combining Eqs. (30) and (34), one obtains the perturbative expression for the fringe velocity,
$v_{x}=\frac{s_{p}}{\hbar k_{R}}\left[G_{ss}-\frac{4E_{R}^{3}N_{v_{x}}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,,$ (36)
with $N_{v_{x}}$ defined in Eq. (68). Both the energy per particle and the chemical potential are even functions of $s_{p}$, reaching their minimum at $s_{p}=0$, where they decrease to the values derived in Ref. [58]. At small $s_{p}$, they increase quadratically [see Figs. 3(a1)-(a2) and 3(b1)-(b2)]. In contrast, both $\Delta\mu$ and $v_{x}$ are odd functions of $s_{p}$, and grow linearly with positive slope for small $|s_{p}|$, as shown in Figs. 3(c1)-(c2) and 3(d1)-(d2). In the limit $s_{p}\to\pm 1$, the energy per particle approaches that of the plane-wave phase [70, 67]. The behavior of $E/N$ as a function of $s_{p}$ depends strongly on the Raman coupling strength. At low $\Omega_{R}$, the quantity $E/N$ increases monotonically with $|s_{p}|$ and attains its maximum at $s_{p}=\pm 1$, as seen in Fig. 3(a1). In contrast, at large $\Omega_{R}$, the quantity $E/N$ develops a global maximum at an intermediate value of $|s_{p}|$ and decreases thereafter, reaching local minima of the energy at $s_{p}=\pm 1$ corresponding to the plane-wave states [see Fig. 3(a2)]. Due to the thermodynamic relation in Eq. (35), this qualitative change also affects the behavior of $\Delta\mu$ and $v_{x}$. While the chemical potential $\mu$ remains a convex function of $s_{p}$, $\Delta\mu$ at large $\Omega_{R}$ develops two stationary extrema, a minimum at negative $s_{p}$ and a maximum at positive $s_{p}$, corresponding to inflection points of the energy per particle. Both $\Delta\mu$ and $v_{x}$ vanish at the local maxima of $E/N$. Interestingly, in the large-$|s_{p}|$ regime, the fringe velocity may acquire the sign opposite to that of the momentum. However, this regime is dynamically unstable, as discussed in Sec. V.2. Finally, within the metastable window between the critical point $\Omega_{\mathrm{cr1}}$ and the spinodal point $\Omega_{\mathrm{sp}}$ the stationary stripe configuration at $s_{p}=0$ represents only a local minimum of the energy per particle, while the plane-wave states at $s_{p}=\pm 1$ are energetically favored. Note that, although $\mu$ and $\Delta\mu$ become ill defined individually at $s_{p}=\pm 1$, the combinations $\mu+\Delta\mu/2$ (for $s_{p}\to 1$) and $\mu-\Delta\mu/2$ (for $s_{p}\to-1$) smoothly approach the chemical potential of the plane-wave phase [67].
V Excitation spectrum of moving stripe patterns
In this section, we examine dynamical properties of moving stripe patterns. We begin by outlining the Bogoliubov theory for studying small oscillations around the stripe background in the comoving frame, where the fringes remain stationary (Sec. V.1). We then compute the Bogoliubov spectrum and the associated sound velocities, highlighting the emergence of both energetic and dynamical instabilities at large momenta (Sec. V.2).
V.1 Bogoliubov theory in the comoving frame
We study the dynamical properties of moving stripe patterns in the comoving frame. Within the framework of Bogoliubov theory [80, 81, 87] small oscillations around the stationary configuration (12) are described by an order parameter of the form
$\Psi^{\prime}(\bm{r}^{\prime},t^{\prime})=\mathrm{e}^{-\mathrm{i}\mu^{\prime}t^{\prime}/\hbar}\mathrm{e}^{-\mathrm{i}m\bm{v}\cdot\bm{r}^{\prime}/\hbar}\left[\Psi_{0}(\bm{r}^{\prime})+\delta\Psi(\bm{r}^{\prime},t^{\prime})\right]\,.$ (37)
To determine the small fluctuation $\delta\Psi(\bm{r}^{\prime},t^{\prime})$, we insert the Ansatz (37) into the Gross-Pitaevskii equation (10). Retaining only terms linear in $\delta\Psi$, we obtain the evolution equation
$\mathrm{i}\hbar\partial_{t^{\prime}}\delta\Psi=\left(h_{\mathrm{SO}}-\mu-\bm{v}\cdot\bm{p}+h_{\mathrm{D}}\right)\delta\Psi+h_{\mathrm{C}}\delta\Psi^{*}\,,$ (38)
where
$\begin{split}h_{\mathrm{D}}={}&{}g_{dd}(\Psi_{0}^{\dagger}\Psi_{0}\mathbb{I}_{2}+\Psi_{0}\Psi_{0}^{\dagger})\\ &{}+g_{ss}\left[(\Psi_{0}^{\dagger}\sigma_{z}\Psi_{0})\sigma_{z}+(\sigma_{z}\Psi_{0})(\sigma_{z}\Psi_{0})^{\dagger}\right]\,,\end{split}$ (39a)
$h_{\mathrm{C}}={}$ ${}g_{dd}\Psi_{0}\Psi_{0}^{T}+g_{ss}(\sigma_{z}\Psi_{0})(\sigma_{z}\Psi_{0})^{T}\,.$ (39b)
As is customary, we look for solutions of the linearized problem (38) with a single, possibly complex, oscillation frequency:
$\delta\Psi(\bm{r}^{\prime},t^{\prime})=U(\bm{r}^{\prime})\mathrm{e}^{-\mathrm{i}\omega t^{\prime}}+V^{*}(\bm{r}^{\prime})\mathrm{e}^{\mathrm{i}\omega^{*}t^{\prime}}\,.$ (40)
The two-component amplitudes $U$ and $V$ depend on position and are defined up to a global complex normalization factor. Substituting Eq. (40) into Eq. (38) and collecting terms oscillating at the same Bogoliubov frequency $\omega$ we obtain an eigenvalue equation for $\omega$ and the corresponding amplitudes:
$(\eta\mathcal{B}-\bm{v}\cdot\bm{p})\begin{pmatrix}U\\ V\end{pmatrix}=\hbar\omega\begin{pmatrix}U\\ V\end{pmatrix}\,,$ (41)
where $\eta=\operatorname{diag}(1,1,-1,-1)$ and
$\mathcal{B}=\begin{pmatrix}h_{\mathrm{SO}}-\mu+h_{\mathrm{D}}&h_{\mathrm{C}}\\ h_{\mathrm{C}}^{*}&(h_{\mathrm{SO}}-\mu+h_{\mathrm{D}})^{*}\end{pmatrix}\,.$ (42)
These expressions formally coincide with those derived for the stationary stripe phase [76, 58], except for the additional Doppler term $-\bm{v}\cdot\bm{p}$ on the left hand side of Eq. (41). As in the $\bm{v}=0$ case, we seek solutions in the form of Bloch waves:
$\begin{pmatrix}U_{b,\bm{k}}(\bm{r}^{\prime})\\ V_{b,\bm{k}}(\bm{r}^{\prime})\end{pmatrix}=\frac{\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{r}^{\prime}}}{\sqrt{V}}\sum_{\text{{\hbox{\bar{m}}} odd}}\begin{pmatrix}\tilde{U}_{b,\bm{k},\bar{m}}\\ \tilde{V}_{b,\bm{k},\bar{m}}\end{pmatrix}\mathrm{e}^{\mathrm{i}\bar{m}k_{1}x^{\prime}}\,.$ (43)
Here, $\bm{k}$ is the excitation quasimomentum, while the $\tilde{U}_{b,\bm{k},\bar{m}}$’s and $\tilde{V}_{b,\bm{k},\bar{m}}$’s are two-component expansion coefficients. Combining Eqs. (41) and (43) we obtain an infinite-dimensional eigenvalue problem for these coefficients and the corresponding Bogoliubov frequencies $\omega_{b,\bm{k}}$. A detailed discussion of the case $\bm{v}=0$ can be found in Ref. [54]. At fixed $\bm{k}$, one obtains an infinite set of eigenfrequencies that form a band structure as $\bm{k}$ varies [76, 58]. This justifies the use of an additional subscript, the band index $b$, to label the amplitudes and the frequency of each Bogoliubov mode. The most general solution to Eq. (38) can then be expressed as a sum over all physically distinct excitation bands and quasimomenta.
The Bogoliubov spectrum obtained by solving the eigenvalue equation (41) exhibits several notable properties:
  1. 1)
    complex frequencies occur in complex conjugate pairs. This follows from the fact that the operator $\eta\mathcal{B}-\bm{v}\cdot\bm{p}$ differs from its Hermitian adjoint by a unitary transformation: $\eta(\eta\mathcal{B}-\bm{v}\cdot\bm{p})\eta^{-1}=(\eta\mathcal{B}-\bm{v}\cdot\bm{p})^{\dagger}$ [87];
  2. 2)
    each mode has a conjugate counterpart that describes the same physical oscillation. Specifically, if the pair $(U_{b,\bm{k}},V_{b,\bm{k}})$ solves Eq. (41) with frequency $\omega_{b,\bm{k}}$, then the pair $(V_{b,\bm{k}}^{*},U_{b,\bm{k}}^{*})$ is a solution with frequency $-\omega_{b,\bm{k}}^{*}$ [87]. However, because of the structure of the Ansatz (40), these two solutions describe the same physical excitation;
  3. 3)
    the spectrum is periodic in the $x$-component of the quasimomentum. Specifically, for each band $b$, the Bogoliubov frequency satisfies the periodicity condition $\omega_{b,\bm{k}+2k_{1}\hat{\bm{e}}_{x}}=\omega_{b,\bm{k}}$, with period $2k_{1}$ corresponding to the extent of the first Brillouin zone.
The first property implies that a moving stripe pattern is dynamically stable as long as its excitation spectrum is entirely real. When this condition is met, the second property ensures that there are equal numbers of solutions of Eq. (41) with positive and negative frequency. These solutions obey the orthonormalization relation [87]
$\begin{split}&{}\int_{V}\mathrm{d}^{3}r^{\prime}\,\left[U_{b^{\prime},\bm{k}^{\prime}}^{\dagger}(\bm{r}^{\prime})U_{b,\bm{k}}(\bm{r}^{\prime})-V_{b^{\prime},\bm{k}^{\prime}}^{\dagger}(\bm{r}^{\prime})V_{b,\bm{k}}(\bm{r}^{\prime})\right]\\ &{}=\mathcal{N}_{b,\bm{k}}\delta_{bb^{\prime}}\delta_{\bm{k}\bm{k^{\prime}}}\,,\end{split}$ (44)
which has two key implications. First, it defines orthogonality between modes belonging to different bands ($b^{\prime}\neq b$) or with different quasimomenta ($\bm{k}^{\prime}\neq\bm{k}$). Second, in the case $b^{\prime}=b$ and $\bm{k}^{\prime}=\bm{k}$, it provides the norm $\mathcal{N}_{b,\bm{k}}$ of a given Bogoliubov mode. Importantly, each physical oscillation corresponds to a pair of solutions with opposite norms. In plotting the Bogoliubov spectrum, such as in Fig. 4, it is customary to include only the frequencies of the positive norm solutions. Furthermore, the system is said to be energetically stable if all positive norm solutions have positive frequency. In the next section, we will show that moving stripe patterns can lose energetic or dynamical stability at sufficiently large momentum. Finally, due to the periodicity of the spectrum, it suffices to consider values of the $x$ component of $\bm{k}$ within the first Brillouin zone. In contrast, the $y$ and $z$ components are unconstrained. Following the conventions of Refs. [76, 58], we define the first Brillouin zone for the excitation quasimomentum as the interval $0\leq k_{x}<2k_{1}$. With this choice, excitations propagating along the positive $x$ direction satisfy $0<k_{x}<k_{1}$, while those along the negative $x$ direction correspond to $k_{1}<k_{x}<2k_{1}$.
V.2 Bogoliubov spectrum and sound velocities
To numerically compute the Bogoliubov spectrum, we truncate the expansions (43) by retaining only a finite number of terms. Specifically, we include components with $-9\leq\bar{m}\leq 9$, following the approach used in Sec. IV for evaluating $\Psi_{0}$. This truncation reduces Eq. (41) to a finite-dimensional eigenvalue problem, which can be readily solved. A detailed discussion of the computational procedure is given in Ref. [54]. In Fig. 4, we display the four lowest-lying bands of the computed spectrum for excitation quasimomentum $\bm{k}$ aligned along the $x$ axis. The figure consists of two rows of panels, corresponding to the two values of the Raman coupling considered in this work. Within each row, we vary the value of $s_{p}$ to illustrate four representative cases:
Figure 4: Excitation spectrum of a spin-orbit-coupled BEC in the stripe phase for various values of $s_{p}$: (a1)-(a2) $s_{p}=0$ (stationary stripes), (b1)-(b2) $s_{p}=0.5$, (c1) $s_{p}=0.95$, (c2) $s_{p}=0.75$, and (d1)-(d2) $s_{p}=0.99$. The top-row panels correspond to $\hbar\Omega_{R}/E_{R}=1.0$, while those in the bottom row correspond to $\hbar\Omega_{R}/E_{R}=2.6$. The excitation quasimomentum is taken along the $x$ direction. Only the four lowest-lying excitation bands are shown; the band index $b$ is indicated next to each curve. In (c1), the inset shows a magnified view near the left edge of the Brillouin zone, where the lowest gapless band exhibits a negative frequency. In all panels, we plot only the real part of the frequency, $\omega_{b,k_{x}}^{\mathrm{R}}$; the imaginary part, $\omega_{b,k_{x}}^{\mathrm{I}}$, is zero except for the lowest gapless band in (c2) (see inset) and in (d2), where it is nonzero but very small and omitted for clarity. The red dashed line in (c2) displays the solution $\omega_{1,\bm{k}}^{*}$ in the dynamically unstable region and $-\omega_{1,2k_{1}\hat{\bm{e}}_{x}-\bm{k}}$ in the dynamically stable region, both obtained from Eq. (41). In panels (d1)-(d2), the black dashed curves represent the two branches of the excitation spectrum in the plane-wave phase, computed using the same parameters as in the case of the stripe-phase spectrum, and with an added Doppler shift (see text for details). The interaction parameters are as in previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
  1. (a)
    Spectrum of stationary stripes ($s_{p}=0$), previously computed in Refs. [76, 58]. The two lowest bands are gapless and exhibit linear dispersion near both edges of the first Brillouin zone, i.e., at $k_{x}=0$ and $k_{x}=2k_{1}$. The two zero-frequency modes at $k_{x}=0$ correspond to the Goldstone modes arising from the spontaneous breaking of global-phase and translation symmetries. Recall that, in the regime $g_{dd}>g_{ss}$ considered here, the upper (lower) phonon branch has predominantly density (spin) character [76], corresponding to the superfluid (crystalline) nature of the stripe phase [58]. The roles are interchanged when $g_{ss}>g_{dd}$.
  2. (b)
    Spectrum of moving stripes at intermediate momentum ($s_{p}=0.5$). Most features remain similar to the stationary case. However, the bands no longer exhibit reflection symmetry with respect to the center of the Brillouin zone, $k_{x}=k_{1}$. This is due to the fact that the transformation $k_{x}\to 2k_{1}-k_{x}$ transforms the spectrum of a stripe pattern to the one with opposite momentum. Additionally, the gaps between bands are narrower for moving stripes, and this trend becomes more pronounced as $|s_{p}|$ increases.
  3. (c)
    Spectrum in the instability regime ($s_{p}=0.95$ for the top row, $s_{p}=0.75$ for the bottom row). The behavior depends on the Raman coupling. At lower Raman coupling (top row), a portion of the lowest gapless band dips into negative frequency, particularly near the left edge of each Brillouin zone (or the right edge for negative $s_{p}$), as shown in the inset of panel (c1). This signals the onset of energetic instability for moving stripes at large momenta. In contrast, at higher Raman coupling (bottom row), the system first exhibits energetic instability (now near the right edge of the Brillouin zone), followed by the emergence of a nonzero imaginary part in the lowest band near both edges of the Brillouin zone, as seen in panel (c2). According to the discussion in the previous section, this indicates that moving stripe patterns become dynamically unstable at high Raman coupling and high momentum.
  4. (d)
    Spectrum near the melting limit ($s_{p}=0.99$). Unlike the previous panels, which focus on the first Brillouin zone, here we plot the spectrum over the range $-4k_{1}\leq k_{x}<2k_{1}$, encompassing three consecutive Brillouin zones. One observes a significant narrowing of the band gaps, heralding their complete closure at the stripe melting point. At this point, segments from different bands and Brillouin zones merge to form the dispersion relations $\omega_{\pm,\bm{k}}-\bm{v}\cdot\bm{k}$ (dashed lines), where $\omega_{\pm,\bm{k}}$ denote the lower (--) and upper ($+$) branches of the excitation spectrum in the plane-wave phase [67]. [^3] Figure 4(d1)-(d2) may suggest that the Bogoliubov frequencies remain periodic in $k_{x}$ even in the plane-wave phase. However, one must recall that $\bm{k}$ represents the excitation quasimomentum. The spectrum loses periodicity when plotted as a function of the excitation momentum, which is well defined only for $|s_{p}|=1$; see the related discussion in Ref. [58]. The Doppler shift $-\bm{v}\cdot\bm{k}$ appears because the spectrum is evaluated in the comoving frame. Notably, the density phonon modes (upper gapless branch) of the stripe phase smoothly evolve into the single phonon branch of the plane-wave phase. More intriguingly, the spin phonon modes (lower gapless branch) transform into the well-known roton minimum of the plane-wave excitation spectrum near $k_{x}=-2k_{1}^{\mathrm{PW}}$ (or $k_{x}=+2k_{1}^{\mathrm{PW}}$ for the plane-wave state with negative condensation momentum, corresponding to $s_{p}=-1$[88, 67, 60, 89, 90].
The occurrence of an energetic instability in the spectrum of moving stripes is consistent with the behavior observed on the plane-wave side of the phase diagram. In this regime, the roton gap progressively decreases as the Raman coupling $\Omega_{R}$ is lowered towards the transition to the stripe phase [67, 89, 90]. Eventually, the gap closes at a spinodal point slightly below the critical value $\Omega_{\mathrm{cr1}}$. Upon further reduction of $\Omega_{R}$, the Bogoliubov modes near the roton minimum acquire negative frequency. This effect can be seen in the dashed lines of Fig. 4(d1)-(d2), although it appears less pronounced in the comoving frame due to the Doppler shift term $-\bm{v}\cdot\bm{k}$. In contrast, the plane-wave spectrum remains dynamically stable for all values of $\Omega_{R}$. This implies that the dynamical instability of moving stripe patterns must vanish once the limit $s_{p}\to\pm 1$ is reached. Our numerical analysis confirms this expectation: beyond a certain threshold, the dynamically unstable region begins to shrink as $|s_{p}|$ increases, and it disappears entirely when $|s_{p}|=1$. This behavior reflects the trend observed in the sound velocities, as discussed below.
Figure 5: Velocities of spin ($b=1$, red lines) and density ($b=2$, blue lines) sound waves as functions of $|s_{p}|$ for (a) $\hbar\Omega_{R}/E_{R}=1.0$ and (b) $\hbar\Omega_{R}/E_{R}=2.6$. Results for both forward ($c_{b,+}$) and backward ($c_{b,-}$) propagation relative to the momentum direction are shown. For the spin sound velocities, we plot both the real parts (solid lines) and the imaginary parts (dashed lines), with the latter being identical for $c_{1,+}$ and $c_{1,-}$. The interaction parameters are the same as in previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
We now analyze in more detail the long wavelength behavior of the two gapless excitation bands ($b=1,2$), focusing in particular on the velocities of sound waves propagating parallel ($c_{b,+}$) and antiparallel ($c_{b,-}$) to the momentum direction. For $s_{p}>0$, these velocities characterize the linear dispersions of the gapless bands near the edges of the first Brillouin zone, namely, $\omega_{b,k_{x}}\simeq c_{b,+}k_{x}$ near $k_{x}=0$ and $\omega_{b,k_{x}}\simeq c_{b,-}(2k_{1}-k_{x})$ near $k_{x}=2k_{1}$. The same relations hold for $s_{p}<0$, with the roles of $c_{b,+}$ and $c_{b,-}$ interchanged. As illustrated in Fig. 5, for both the spin ($b=1$) and density ($b=2$) sound modes, the velocities in the direction of the momentum differ from those in the opposite direction, except at $s_{p}=0$, where the two become equal. This symmetry at $s_{p}=0$ reflects the underlying symmetry of the excitation spectrum with respect to the center of the Brillouin zone, as discussed earlier.
The density sound velocities $c_{2,\pm}$ are always real and positive. In the limit $|s_{p}|\to 1$, they approach the expressions $c_{2,\pm}\to c_{\pm x}\mp v_{x}$, where $c_{+x}$ ($c_{-x}$) denotes the speed of density sound waves propagating along (against) the $x$ direction in the plane-wave phase. Note that $c_{+x}\neq c_{-x}$ in the presence of nonzero spin-dependent interactions ($g_{ss}\neq 0$[67].
In contrast, the behavior of the spin sound velocities reveals the onset of instability. As $|s_{p}|$ increases, one of the two velocities $c_{1,\pm}$ becomes negative at a certain point, indicating the emergence of an energetic instability, as previously discussed. For high Raman coupling, this instability is soon followed by the appearance of complex sound velocities: both $c_{1,+}$ and $c_{1,-}$ acquire imaginary parts of equal magnitude. These imaginary parts first grow with $|s_{p}|$, then decrease, and vanish as $|s_{p}|\to 1$. In this limit, the real parts tend towards opposite values, satisfying $c_{1,+}=-c_{1,-}$. This behavior confirms that the emergence of energetic instability is a necessary precursor to ensuring continuity of the excitation spectrum slope in the comoving frame at the points $k_{x}=\pm 2k_{1}^{\mathrm{PW}}$ in the plane-wave phase.
VI Conclusions
In this work, we have investigated the properties of traveling supersolid patterns in a spin-orbit-coupled Bose-Einstein condensate. The motion of the density fringes stems from a population imbalance between the two minima of the single-particle dispersion. Antiferromagnetic spin-dependent interactions create a chemical potential difference which sets the stripe velocity. These moving stripes exhibit asymmetric densities of the two spin components, notable spin polarization, reduced contrast, and higher energy than that of stationary stripes. As momentum increases, these effects intensify, eventually leading to the disappearance of the fringes and the emergence of a uniform plane-wave condensate. This indicates that traveling stripes interpolate between the stationary stripe and plane-wave phases. By analyzing the excitation spectrum, we uncover energetic and dynamical instabilities at high momentum, mainly affecting the spin phonon branch that becomes the mode with the roton minimum in the plane-wave phase. Our findings extend prior work [65, 66] on the slow motion of stripes induced by the release of a weak spin perturbation, where the propagation of fringes manifests the crystal Goldstone mode of the stripe phase.
The results presented in this work may be relevant for ongoing experimental investigations of spin-orbit-coupled Bose gases. Notably, moving stripe patterns have already been observed by Ketterle’s group at MIT [20]. In their experiment, a Raman detuning was employed to generate a chemical potential difference between the two minima of the single-particle spectrum, thereby inducing stripe motion without altering the atomic population balance. In contrast, the configuration explored in this work involves population imbalance between the two momentum components, with the resulting stripe motion driven by spin-dependent interactions. This scenario can be realized experimentally by first preparing the Bose-Einstein condensate in a superposition of two atomic hyperfine states with different weights. Raman coupling can then be adiabatically introduced and ramped up to the desired strength. As in Ref. [20], the resulting moving density modulations could be detected using Bragg scattering techniques.
Acknowledgements.
Useful discussions and collaborations with K. T. Geier, Ph. Hauke, W. Ketterle, Y. Li, A. Recati, S. Stringari, and D. Trypogeorgos are acknowledged. This work was supported by the Italian Ministry of University and Research (MUR) through the PNRR MUR project: ‘National Quantum Science and Technology Institute’ - NQSTI (PE0000023) and the PNRR MUR project: ‘Integrated Infrastructure Initiative in Photonic and Quantum Sciences’ - I-PHOQS (IR0000016). We acknowledge the support of the Quantum Optical Networks based on Exciton-polaritons - (Q-ONE) funding from the HORIZON-EIC-2022-PATHFINDER CHALLENGES EU programme under grant agreement No. 101115575, and of the Neuromorphic Polariton Accelerator - (PolArt) funding from the Horizon-EIC-2023-Pathfinder Open EU programme under grant agreement No. 101130304. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Innovation Council and SMEs Executive Agency (EISMEA). Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Perturbative method
This appendix presents a detailed methodology employed to derive the perturbative formulas discussed in the main text. The general procedure described in Sec. A.1 closely follows the approach of Ref. [58], with appropriate generalizations to incorporate the additional momentum constraint. The first and second order corrections to the order parameter are subsequently derived in Sec. A.2 and Sec. A.3, respectively.
A.1 General outline of the perturbative approach
To perturbatively solve the time-independent Gross-Pitaevskii equation with a momentum constraint, Eq. (8), we first observe that the stripe order parameter (14) depends solely on the combination $k_{1}x^{\prime}$. This allows us to neglect any derivatives with respect to $y^{\prime}$ and $z^{\prime}$, and to assume that $\Psi_{0}$ is a function of the single dimensionless variable $X=k_{1}x^{\prime}$. This simplification is especially useful because the optimal wave vector $k_{1}$ depends on the Raman coupling $\Omega_{R}$, which in turn can cause the antiperiod of $\Psi_{0}$, $\pi/k_{1}$, to vary at each perturbative order (although, by symmetry, only even order corrections in $\Omega_{R}$ are permitted; see Sec. IV.3). After this change of variable, the antiperiodicity condition becomes $\Psi_{0}(X+\pi)=-\Psi_{0}(X)$, fixing the antiperiod at $\pi$ for all orders. Therefore, $\Psi_{0}$ is $2\pi$-periodic in $X$.
With all the above considerations in place, we now insert the perturbative expansions [Eqs. (21), (22), (23), and (24)] into Eq. (8) and collect terms of equal order. At perturbative order $l\geq 1$, we obtain the recurrence relation
$\begin{split}&{}\left[E_{R}\left(-\mathrm{i}\nabla_{X}-\sigma_{z}\right)^{2}-G_{ss}s_{p}\left(-\mathrm{i}\nabla_{X}-\sigma_{z}\right)+\mathcal{L}_{D}\right]\Psi_{0}^{(l)}\\ &{}+\mathcal{L}_{C}\Psi_{0}^{(l)*}=\left[\mu^{(l)}+\frac{\Delta\mu^{(l)}}{2}(-\mathrm{i}\nabla_{X})\right]\Psi_{0}^{(0)}-\mathcal{J}^{(l)}\,,\end{split}$ (45)
where $\nabla_{X}$ denotes derivation with respect to the dimensionless variable $X$. The matrices $\mathcal{L}_{\mathrm{D}}$ and $\mathcal{L}_{\mathrm{C}}$ are defined as
$\mathcal{L}_{\mathrm{D}}$ ${}=\frac{1}{2}\begin{pmatrix}(1+s_{p})(G_{dd}+G_{ss})&\sqrt{1-s_{p}^{2}}(G_{dd}-G_{ss})\mathrm{e}^{2\mathrm{i}X}\\ \sqrt{1-s_{p}^{2}}(G_{dd}-G_{ss})\mathrm{e}^{-2\mathrm{i}X}&(1-s_{p})(G_{dd}+G_{ss})\end{pmatrix}\,,$ (46a)
$\mathcal{L}_{\mathrm{C}}$ ${}=\frac{1}{2}\begin{pmatrix}(1+s_{p})(G_{dd}+G_{ss})\mathrm{e}^{2\mathrm{i}X}&\sqrt{1-s_{p}^{2}}(G_{dd}-G_{ss})\\ \sqrt{1-s_{p}^{2}}(G_{dd}-G_{ss})&(1-s_{p})(G_{dd}+G_{ss})\mathrm{e}^{-2\mathrm{i}X}\end{pmatrix}\,,$ (46b)
and the source term $\mathcal{J}^{(l)}$ reads
$\mathcal{J}^{(l)}={}$ ${}-\frac{\hbar^{2}}{2m}\sum_{l_{2}=2}^{l}\sum_{l_{1}=1}^{l_{2}-1}k_{1}^{(l_{1})}k_{1}^{(l_{2}-l_{1})}\nabla_{X}^{2}\Psi_{0}^{(l-l_{2})}$
${}+\frac{\hbar^{2}k_{R}}{m}\sum_{l_{1}=1}^{l-1}k_{1}^{(l_{1})}\left(-\nabla_{X}^{2}+\mathrm{i}\sigma_{z}\nabla_{X}\right)\Psi_{0}^{(l-l_{1})}$
${}+\frac{\hbar\Omega_{R}}{2}\,\sigma_{x}\Psi_{0}^{(l-1)}$
${}-\sum_{l_{1}=1}^{l-1}\mu^{(l_{1})}\Psi_{0}^{(l-l_{1})}-\sum_{l_{1}=1}^{l-1}\frac{\Delta\mu^{(l_{1})}}{2}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l-l_{1})}$
${}+g_{dd}\sum_{l_{1},l_{2},l_{3}=0}^{l-1}\left[\Psi_{0}^{(l_{1})\dagger}\Psi_{0}^{(l_{2})}\right]\Psi_{0}^{(l_{3})}\delta_{l,l_{1}+l_{2}+l_{3}}$
${}+g_{ss}\sum_{l_{1},l_{2},l_{3}=0}^{l-1}\left[\Psi_{0}^{(l_{1})\dagger}\sigma_{z}\Psi_{0}^{(l_{2})}\right]\sigma_{z}\Psi_{0}^{(l_{3})}\delta_{l,l_{1}+l_{2}+l_{3}}\,.$ (47)
In deriving Eqs. (46), we have set $\theta=\Delta\theta=0$ in the zero coupling expression (18) for the stripe order parameter, without loss of generality. Equation (45) thus provides a recursive method for determining the $l$th order corrections to the order parameter $\Psi_{0}^{(l)}$, as well as to the average chemical potential $\mu^{(l)}$ and the chemical potential difference $\Delta\mu^{(l)}$, once all lower-order corrections [which enter the source term $\mathcal{J}^{(l)}$] are known. As in standard time-independent perturbation theory, the first step is to extract $\mu^{(l)}$ and $\Delta\mu^{(l)}$. To do this, we multiply both sides of Eq. (45) by $\Psi_{0}^{(0)\dagger}$ and integrate over the full period of the wave function (from $X=-\pi$ to $X=\pi$). We then repeat this procedure with the operator $\Psi_{0}^{(0)\dagger}(-\mathrm{i}\nabla_{X})$. This yields two linear equations for $\mu^{(l)}$ and $\Delta\mu^{(l)}$, that can be readily solved. While the resulting expressions typically depend on the yet unknown $\Psi_{0}^{(l)}$, this dependence appears only through two integrals, which can be expressed using lower order quantities. These are
$\begin{split}&{}\int_{-\pi}^{\pi}\mathrm{d}X\left[\Psi_{0}^{(0)\dagger}\Psi_{0}^{(l)}+\Psi_{0}^{(l)\dagger}\Psi_{0}^{(0)}\right]\\ &{}=-\sum_{l_{1}=1}^{l-1}\int_{-\pi}^{\pi}\mathrm{d}X\,\Psi_{0}^{(l_{1})\dagger}\Psi_{0}^{(l-l_{1})}\end{split}$ (48)
and
$\begin{split}&{}\int_{-\pi}^{\pi}\mathrm{d}X\left[\Psi_{0}^{(0)\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l)}+\Psi_{0}^{(l)\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(0)}\right]\\ &{}=-\sum_{l_{1}=1}^{l-1}\int_{-\pi}^{\pi}\mathrm{d}X\,\Psi_{0}^{(l_{1})\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l-l_{1})}\,.\end{split}$ (49)
These two relations are obtained by imposing the conservation of particle number [Eq. (4)] and momentum [Eq. (5)] at each perturbative order.
The general solution of the nonhomogeneous linear equation (45) can be constructed by adding a particular solution to the general solution $\Psi_{\mathrm{Hom}}^{(l)}$ of the associated homogeneous equation. The latter is a linear combination of eight linearly independent functions with arbitrary coefficients. As in the $s_{p}=0$ case discussed in Ref. [58], six of these functions violate the periodicity condition of the order parameter and must therefore be excluded. This leaves only two admissible functions, and $\Psi_{\mathrm{Hom}}^{(l)}$ reduces to the form
$\Psi_{\mathrm{Hom}}^{(l)}=\mathrm{i}\left[\theta^{(l)}+\Delta\theta^{(l)}(-\mathrm{i}\nabla_{X})\right]\Psi_{0}^{(0)}\,,$ (50)
where $\theta^{(l)}$ and $\Delta\theta^{(l)}$ are undetermined constants. To maintain consistency with the perturbative expansion, both coefficients must scale as $(\hbar\Omega_{R}/4E_{R})^{l}$. The two terms in Eq. (50) correspond to infinitesimal shifts in the condensate phase and in the position of density fringes, respectively, which reflects the spontaneous breaking of global $\mathrm{U}(1)$ and translation symmetry in the stripe phase. To uniquely determine $\theta^{(l)}$ and $\Delta\theta^{(l)}$, we impose the following orthogonality conditions on the order-$l$ correction to the order parameter:
${}\int_{-\pi}^{\pi}\mathrm{d}X\,\operatorname{Im}\left[\Psi_{0}^{(0)\dagger}\Psi_{0}^{(l)}\right]=0\,,$ (51a)
${}\int_{-\pi}^{\pi}\mathrm{d}X\,\operatorname{Im}\left[\Psi_{0}^{(0)\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l)}\right]=0\,.$ (51b)
The function $\Psi_{0}^{(l)}$ computed following the above prescriptions is a sum of plane-wave components whose wave vectors (in units of $k_{1}$) are all odd integers from $-(2l+1)$ to $12l+$ 1. Thus, the number of harmonics increases with the perturbative order, in agreement with the Bloch-wave structure of the exact order parameter, Eq. (14). In the next sections, we explicitly compute the corrections at first and second order in $\hbar\Omega_{R}/4E_{R}$.
A.2 First order results
For $l=1$, the source term (47) receives contributions solely from the Raman coupling, yielding
$\mathcal{J}^{(1)}=\sqrt{\bar{n}}\left[\begin{pmatrix}0\\ \mathcal{J}_{+1,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}\mathcal{J}_{-1,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\right]\,,$ (52)
with
$\mathcal{J}_{+1,\downarrow}^{(1)}(s_{p})=\mathcal{J}_{-1,\uparrow}^{(1)}(-s_{p})=\sqrt{\frac{1+s_{p}}{2}}\frac{\hbar\Omega_{R}}{2}\,.$ (53)
Using this expression and following the procedure outlined in Sec. A.1, one immediately finds that $\mu^{(1)}=\Delta\mu^{(1)}=0$. Furthermore, the first order correction to the order parameter, satisfying both Eq. (45) and the orthogonality conditions (51), takes the form
$\begin{split}\Psi_{0}^{(1)}={}&{}\sqrt{\bar{n}}\bigg{[}\begin{pmatrix}\tilde{\Psi}_{+3,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{+1,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{-1,\uparrow}^{(1)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-3,\downarrow}^{(1)}(s_{p})\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}\bigg{]}\,,\end{split}$ (54)
where the coefficients are given by
$\tilde{\Psi}_{+1,\downarrow}^{(1)}(s_{p})$ ${}=\tilde{\Psi}_{-1,\uparrow}^{(1)}(-s_{p})=\sqrt{\frac{1+s_{p}}{2}}\frac{N_{1}^{(1)}}{2D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,,$ (55a)
$\begin{split}\tilde{\Psi}_{+3,\uparrow}^{(1)}(s_{p})&{}=\tilde{\Psi}_{-3,\downarrow}^{(1)}(-s_{p})\\ &{}=\sqrt{\frac{1-s_{p}}{2}}\frac{(1+s_{p})N_{3}^{(1)}}{2D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,.\end{split}$ (55b)
The explicit forms of $D_{1}$, $D_{2}$, $N_{1}^{(1)}$, and $N_{3}^{(1)}$ are given in Appendix B. These results are obtained by inserting the Ansatz (54) into Eq. (45), grouping terms with the same oscillatory behavior, and equating coefficients on both sides of the equation. Once $\Psi_{0}^{(1)}$ is known, any observable can be computed to first order in $\hbar\Omega_{R}/4E_{R}$. Among those discussed in Sec. IV.3, only the density profiles [see Eqs. (25), (26), and (27)] and the associated contrasts exhibit nonzero first order corrections.
A.3 Second order results
At second order in the Raman coupling ($l=2$), the source term takes the form
$\begin{split}\mathcal{J}^{(2)}={}&{}\sqrt{\bar{n}}\bigg{[}\begin{pmatrix}\mathcal{J}_{+5,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{5\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{+3,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\mathcal{J}_{+1,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{-1,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\mathcal{J}_{-3,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}+\begin{pmatrix}0\\ \mathcal{J}_{-5,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-5\mathrm{i}X}\bigg{]}\,.\end{split}$ (56)
The coefficients in this expression can be derived from the general formula (47), using the results of the previous section. They satisfy the symmetry relations $\mathcal{J}_{+5,\uparrow}^{(2)}(s_{p})=\mathcal{J}_{-5,\downarrow}^{(2)}(-s_{p})$, $\mathcal{J}_{+3,\downarrow}^{(2)}(s_{p})=\mathcal{J}_{-3,\uparrow}^{(2)}(-s_{p})$, and $\mathcal{J}_{+1,\uparrow}^{(2)}(s_{p})=\mathcal{J}_{-1,\downarrow}^{(2)}(-s_{p})$. Applying the method of Sec. A.1, one obtains the second order corrections to the chemical potentials $\mu^{(2)}$ and $\Delta\mu^{(2)}$, which are reported in the main text [see second term on the right hand side of Eq. (33) and Eq. (34), respectively]. The second order correction to the order parameter, derived in the same manner as the first order one (see Sec. A.2), reads
$\begin{split}\Psi_{0}^{(2)}={}&{}\sqrt{\bar{n}}\bigg{[}\begin{pmatrix}\tilde{\Psi}_{+5,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{5\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{+3,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{3\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{+1,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-1,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-\mathrm{i}X}\\ {}&{}+\begin{pmatrix}\tilde{\Psi}_{-3,\uparrow}^{(2)}(s_{p})\\ 0\end{pmatrix}\mathrm{e}^{-3\mathrm{i}X}+\begin{pmatrix}0\\ \tilde{\Psi}_{-5,\downarrow}^{(2)}(s_{p})\end{pmatrix}\mathrm{e}^{-5\mathrm{i}X}\bigg{]}\,,\end{split}$ (57)
with the coefficients given by
$\begin{split}\tilde{\Psi}_{+1,\uparrow}^{(2)}(s_{p})&{}=\tilde{\Psi}_{-1,\downarrow}^{(2)}(-s_{p})\\ &{}=\sqrt{\frac{1+s_{p}}{2}}\frac{N_{1}^{(2)}}{4D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,\end{split}$ (58a)
$\begin{split}\tilde{\Psi}_{+3,\downarrow}^{(2)}(s_{p})&{}=\tilde{\Psi}_{-3,\uparrow}^{(2)}(-s_{p})\\ &{}=-\sqrt{\frac{1-s_{p}}{2}}\frac{(1+s_{p})N_{3}^{(2)}}{4D_{1}^{2}D_{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,\end{split}$ (58b)
$\begin{split}\tilde{\Psi}_{+5,\uparrow}^{(2)}(s_{p})&{}=\tilde{\Psi}_{-5,\downarrow}^{(2)}(-s_{p})\\ &{}=\sqrt{\frac{1+s_{p}}{2}}\frac{(1-s_{p}^{2})N_{5}^{(2)}}{4D_{1}^{2}D_{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,.\end{split}$ (58c)
The quantities $D_{1}$, $D_{2}$, $N_{1}^{(2)}$, $N_{3}^{(2)}$, and $N_{5}^{(2)}$ are given in Appendix B. The structure of Eqs. (58) respects the orthogonality constraints (51). With the expression for $\Psi_{0}^{(2)}$ at hand, the second order corrections to observables can be readily computed. Selected results are presented in Sec. IV.3 of the main text. Finally, it is worth noting that all the perturbative expressions reported in this appendix reduce to those of Ref. [58] when $s_{p}=0$.
Appendix B Coefficients of perturbative formulas
In this Appendix we give the explicit forms of the coefficients introduced in Secs. IV.3, A.2, and A.3. These coefficients appear in various perturbative formulas.
The two quantities appearing in the denominators of all the perturbative expressions are
$D_{1}={}$ ${}4E_{R}^{2}\left(2E_{R}+G_{ss}\right)\left(2E_{R}+G_{dd}\right)-2E_{R}G_{ss}$
×${}\times\left[\left(4E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}^{2}$
${}+G_{ss}^{4}s_{p}^{4}$ (59)
and
$D_{2}={}$ ${}4E_{R}^{2}\left(8E_{R}+G_{ss}\right)\left(8E_{R}+G_{dd}\right)-2E_{R}G_{ss}$
×${}\times\left[\left(16E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}^{2}$
${}+G_{ss}^{4}s_{p}^{4}\,.$ (60)
The coefficients entering the numerators of the observables computed in Sec. IV.3 are
$N_{\tilde{n}_{2}}={}$ ${}48E_{R}^{5}(2E_{R}+G_{ss})^{2}(8E_{R}+G_{ss})(4E_{R}+G_{dd})G_{dd}$
${}+16E_{R}^{4}G_{ss}\big{[}E_{R}\left(16E_{R}^{2}-16E_{R}G_{ss}-3G_{ss}^{2}\right)G_{ss}^{2}$
${}+\left(4E_{R}^{3}-48E_{R}^{2}G_{ss}-31E_{R}G_{ss}^{2}-3G_{ss}^{3}\right)G_{ss}G_{dd}$
${}+\left(12E_{R}^{3}+8E_{R}^{2}G_{ss}-E_{R}G_{ss}^{2}-G_{ss}^{3}\right)G_{dd}^{2}\big{]}s_{p}^{2}$
${}-4E_{R}^{3}G_{ss}^{3}\Big{[}\left(24E_{R}^{2}-20E_{R}G_{ss}-3G_{ss}^{2}\right)G_{ss}^{2}$
${}+4\left(6E_{R}^{2}+E_{R}G_{ss}-G_{ss}^{2}\right)G_{ss}G_{dd}$
${}+\left(2E_{R}+G_{ss}\right)^{2}G_{dd}^{2}\Big{]}s_{p}^{4}$
${}+4E_{R}^{2}G_{ss}^{6}[\left(3E_{R}-G_{ss}\right)G_{ss}$
${}+\left(2E_{R}+G_{ss}\right)G_{dd}]s_{p}^{6}-E_{R}G_{ss}^{9}s_{p}^{8}\,,$ (61)
$N_{\tilde{s}_{z,2}}={}$ ${}16E_{R}^{4}(2E_{R}+G_{ss})\big{[}128E_{R}^{3}G_{ss}^{2}$
${}+32E_{R}^{2}\left(7E_{R}+3G_{ss}\right)G_{ss}G_{dd}$
${}+2E_{R}\left(72E_{R}^{2}+58E_{R}G_{ss}+9G_{ss}^{2}\right)G_{dd}^{2}$
${}+\left(12E_{R}+G_{ss}\right)\left(2E_{R}+G_{ss}\right)G_{dd}^{3}\big{]}$
${}-8E_{R}^{3}G_{ss}^{2}\Big{[}32E_{R}^{2}\left(7E_{R}+3G_{ss}\right)G_{ss}^{2}$
${}+4E_{R}\left(62E_{R}^{2}+53E_{R}G_{ss}+9G_{ss}^{2}\right)G_{ss}G_{dd}$
${}+\left(56E_{R}^{3}+108E_{R}^{2}G_{ss}+46E_{R}G_{ss}^{2}+3G_{ss}^{3}\right)G_{dd}^{2}$
${}+\left(2E_{R}+G_{ss}\right)^{2}G_{dd}^{3}\Big{]}s_{p}^{2}$
${}+4E_{R}^{2}G_{ss}^{4}\big{[}18E_{R}\left(4E_{R}+G_{ss}\right)G_{ss}^{2}$
${}+\left(56E_{R}^{2}+44E_{R}G_{ss}+3G_{ss}^{2}\right)G_{ss}G_{dd}$
${}+\left(4E_{R}^{2}+8E_{R}G_{ss}+3G_{ss}^{2}\right)G_{dd}^{2}\big{]}s_{p}^{4}$
${}-2E_{R}G_{ss}^{7}[\left(14E_{R}+G_{ss}\right)G_{ss}$
${}+\left(4E_{R}+3G_{ss}\right)G_{dd}]s_{p}^{6}+G_{ss}^{10}s_{p}^{8}\,,$ (62)
$N_{k_{1}}={}$ ${}4E_{R}^{2}\left(2E_{R}+G_{ss}\right)^{2}\left(4E_{R}^{2}+2E_{R}G_{dd}+G_{dd}^{2}\right)$
${}+4E_{R}\big{[}4E_{R}^{3}\left(4E_{R}+3G_{ss}\right)G_{ss}+\big{(}8E_{R}^{4}$
${}+16E_{R}^{3}G_{ss}+4E_{R}^{2}G_{ss}^{2}-3E_{R}G_{ss}^{3}-G_{ss}^{4}\big{)}G_{dd}\big{]}s_{p}^{2}$
${}-G_{ss}^{2}\big{[}\left(16E_{R}^{3}-16E_{R}^{2}G_{ss}+2E_{R}G_{ss}^{2}+G_{ss}^{3}\right)G_{ss}$
${}-4E_{R}^{2}\left(2E_{R}+G_{ss}\right)G_{dd}\big{]}s_{p}^{4}-2E_{R}G_{ss}^{5}s_{p}^{6}\,,$ (63)
$N_{\langle\sigma_{z}\rangle}={}$ ${}4E_{R}^{2}\big{[}2E_{R}\left(4E_{R}^{2}+6G_{ss}E_{R}+G_{ss}^{2}\right)$
${}+2E_{R}\left(4E_{R}+3G_{ss}\right)G_{dd}+\left(2E_{R}+G_{ss}\right)G_{dd}^{2}\big{]}$
${}-4E_{R}G_{ss}\left[3E_{R}G_{ss}^{2}-\left(2E_{R}^{2}-G_{ss}^{2}\right)G_{dd}\right]s_{p}^{2}$
${}-(2E_{R}-G_{ss})G_{ss}^{4}s_{p}^{4}\,,$ (64)
$N_{E}={}$ ${}2E_{R}\left(2E_{R}+G_{ss}\right)\left(4E_{R}+G_{dd}\right)$
${}+\left[\left(8E_{R}^{2}-G_{ss}^{2}\right)G_{ss}+2E_{R}\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}^{2}$
${}-G_{ss}^{3}s_{p}^{4}\,,$ (65)
$N_{\mu}={}$ ${}-4E_{R}^{3}\left(2E_{R}+G_{ss}\right)^{2}\left(8E_{R}^{2}+4E_{R}G_{dd}+G_{dd}^{2}\right)$
${}+4E_{R}^{2}\big{[}E_{R}\left(24E_{R}^{2}+20E_{R}G_{ss}+3G_{ss}^{2}\right)G_{ss}^{2}$
${}+\left(16E_{R}^{3}+24E_{R}^{2}G_{ss}+10E_{R}G_{ss}^{2}+G_{ss}^{3}\right)G_{ss}G_{dd}$
${}+\left(4E_{R}^{3}+8E_{R}^{2}G_{ss}+5E_{R}G_{ss}^{2}+G_{ss}^{3}\right)G_{dd}^{2}\big{]}s_{p}^{2}$
${}-E_{R}G_{ss}^{3}\big{[}\left(20E_{R}^{2}+16E_{R}G_{ss}+G_{ss}^{2}\right)G_{ss}$
${}+\left(8E_{R}^{2}+12E_{R}G_{ss}+4G_{ss}^{2}\right)G_{dd}\big{]}s_{p}^{4}$
${}+\left(E_{R}+G_{ss}\right)G_{ss}^{6}s_{p}^{6}\,,$ (66)
$N_{\Delta\mu}={}$ ${}4E_{R}^{2}\Big{[}2E_{R}\left(8E_{R}^{2}+8E_{R}G_{ss}+G_{ss}^{2}\right)G_{ss}$
${}+4E_{R}\left(2E_{R}^{2}+5E_{R}G_{ss}+2G_{ss}^{2}\right)G_{dd}$
${}+\left(2E_{R}+G_{ss}\right)^{2}G_{dd}^{2}\Big{]}$
${}-4E_{R}G_{ss}^{3}[4E_{R}\left(E_{R}+G_{ss}\right)$
${}+\left(2E_{R}+G_{ss}\right)G_{dd}]s_{p}^{2}+G_{ss}^{6}s_{p}^{4}\,,$ (67)
and
$N_{v_{x}}={}$ ${}4E_{R}^{2}(2E_{R}+G_{ss})\big{[}\left(4E_{R}^{2}+6E_{R}G_{ss}+G_{ss}^{2}\right)G_{ss}$
${}+\left(4E_{R}^{2}+8E_{R}G_{ss}+3G_{ss}^{2}\right)G_{dd}$
${}+\left(2E_{R}+G_{ss}\right)G_{dd}^{2}\big{]}-2E_{R}G_{ss}\big{[}4\big{(}4E_{R}^{3}$
${}+5E_{R}^{2}G_{ss}+3E_{R}G_{ss}^{2}+G_{ss}^{3}\big{)}G_{ss}+\big{(}8E_{R}^{3}$
${}+16E_{R}^{2}G_{ss}+8E_{R}G_{ss}^{2}+G_{ss}^{3}\big{)}G_{dd}\big{]}s_{p}^{2}$
${}+2E_{R}G_{ss}^{3}[4\left(E_{R}+G_{ss}\right)G_{ss}$
${}-\left(2E_{R}+G_{ss}\right)G_{dd}]s_{p}^{4}+G_{ss}^{6}s_{p}^{6}\,.$ (68)
In the numerators of the first order corrections to the Fourier expansion coefficients of the order parameter, evaluated in Appendix A.2, one has
$N_{1}^{(1)}={}$ ${}-2E_{R}^{2}\left(4E_{R}+G_{dd}\right)\left(2E_{R}+G_{ss}\right)-2E_{R}^{2}$
×${}\times\left[\left(4E_{R}+G_{ss}\right)G_{ss}+\left(2E_{R}+G_{ss}\right)G_{dd}\right]s_{p}$
${}+E_{R}\left(2E_{R}+G_{ss}\right)G_{ss}^{2}s_{p}^{2}+E_{R}G_{ss}^{3}s_{p}^{3}$ (69)
and
$N_{3}^{(1)}={}$ ${}2E_{R}^{2}\left(2E_{R}+G_{ss}\right)G_{dd}$
${}-2E_{R}^{2}G_{ss}^{2}s_{p}-E_{R}G_{ss}^{3}s_{p}^{2}\,.$ (70)
Finally, the second order corrections to the Fourier expansion coefficients of the order parameter (see Appendix A.3) depend on the following quantities:
$N_{1}^{(2)}={}$ ${}-4E_{R}^{4}(2E_{R}+G_{ss})^{2}\left(8E_{R}^{2}+4E_{R}G_{dd}+G_{dd}^{2}\right)$
${}-8E_{R}^{4}(2E_{R}+G_{ss})\big{[}2E_{R}\left(4E_{R}+G_{ss}\right)G_{ss}$
${}+\left(4E_{R}^{2}+4E_{R}G_{ss}-G_{ss}^{2}\right)G_{dd}$
${}+\left(2E_{R}+G_{ss}\right)G_{dd}^{2}\big{]}s_{p}$
${}+4E_{R}^{3}\Big{[}E_{R}\left(4E_{R}+G_{ss}\right)G_{ss}^{3}-\big{(}8E_{R}^{3}+8E_{R}^{2}G_{ss}$
${}-G_{ss}^{3}\big{)}G_{ss}G_{dd}+E_{R}\left(2E_{R}+G_{ss}\right)^{2}G_{dd}^{2}\Big{]}s_{p}^{2}$
${}+4E_{R}^{3}G_{ss}^{3}\big{[}8E_{R}^{2}+4E_{R}G_{ss}-G_{ss}^{2}$
${}+2\left(2E_{R}+G_{ss}\right)G_{dd}\big{]}s_{p}^{3}$
${}-E_{R}^{2}G_{ss}^{3}(2E_{R}+G_{ss})\big{(}G_{ss}^{2}-6E_{R}G_{ss}$
${}+4E_{R}G_{dd}\big{)}s_{p}^{4}-2E_{R}^{2}G_{ss}^{6}s_{p}^{5}+E_{R}^{2}G_{ss}^{6}s_{p}^{6}\,,$ (71)
$N_{3}^{(2)}={}$ ${}16E_{R}^{7}\left(2E_{R}+G_{ss}\right)^{2}\left(8E_{R}+G_{ss}\right)(16E_{R}$
${}+5G_{dd})G_{dd}-8E_{R}^{6}(2E_{R}+G_{ss})\big{[}32E_{R}^{2}G_{ss}^{3}$
${}-8E_{R}\left(48E_{R}^{2}+15E_{R}G_{ss}-G_{ss}^{2}\right)G_{ss}G_{dd}$
${}-\left(160E_{R}^{3}+168E_{R}^{2}G_{ss}+38E_{R}G_{ss}^{2}+G_{ss}^{3}\right)G_{dd}^{2}$
${}-\left(16E_{R}+G_{ss}\right)\left(2E_{R}+G_{ss}\right)G_{dd}^{3}\big{]}s_{p}$
${}+8E_{R}^{6}G_{ss}\Big{[}4E_{R}\left(16E_{R}^{2}-4E_{R}G_{ss}-3G_{ss}^{2}\right)G_{ss}^{2}$
${}+\big{(}48E_{R}^{3}-92E_{R}^{2}G_{ss}-74E_{R}G_{ss}^{2}$
${}-11G_{ss}^{3}\big{)}G_{ss}G_{dd}+2E_{R}\big{(}36E_{R}^{2}+28E_{R}G_{ss}$
${}+5G_{ss}^{2}\big{)}G_{dd}^{2}+2\left(2E_{R}+G_{ss}\right)^{2}G_{dd}^{3}\Big{]}s_{p}^{2}$
${}+4E_{R}^{5}G_{ss}\big{[}2E_{R}\left(48E_{R}^{2}+12E_{R}G_{ss}+7G_{ss}^{2}\right)G_{ss}^{3}$
${}-2\big{(}44E_{R}^{3}+80E_{R}^{2}G_{ss}+26E_{R}G_{ss}^{2}$
${}+G_{ss}^{3}\big{)}G_{ss}^{2}G_{dd}-\big{(}8E_{R}^{3}+76E_{R}^{2}G_{ss}+42E_{R}G_{ss}^{2}$
${}+3G_{ss}^{3}\big{)}G_{ss}G_{dd}^{2}+\big{(}8E_{R}^{3}+4E_{R}^{2}G_{ss}-2E_{R}G_{ss}^{2}$
${}-G_{ss}^{3}\big{)}G_{dd}^{3}\big{]}s_{p}^{3}+4E_{R}^{5}G_{ss}^{2}\Big{[}2\left(5E_{R}+3G_{ss}\right)G_{ss}^{4}$
${}+\left(20E_{R}^{2}+16E_{R}G_{ss}+G_{ss}^{2}\right)G_{ss}^{2}G_{dd}$
${}+\left(4E_{R}^{2}-4E_{R}G_{ss}-3G_{ss}^{2}\right)G_{ss}G_{dd}^{2}$
${}+\left(2E_{R}+G_{ss}\right)^{2}G_{dd}^{3}\Big{]}s_{p}^{4}$
${}-2E_{R}^{4}G_{ss}^{4}\big{[}\left(48E_{R}^{2}-6E_{R}G_{ss}-G_{ss}^{2}\right)G_{ss}^{2}$
${}-\left(4E_{R}^{2}+22E_{R}G_{ss}+3G_{ss}^{2}\right)G_{ss}G_{dd}$
${}+2\left(4E_{R}^{2}-G_{ss}^{2}\right)G_{dd}^{2}\big{]}s_{p}^{5}$
${}-2E_{R}^{4}G_{ss}^{5}\big{[}(18E_{R}+G_{ss})G_{ss}^{2}$
${}+4E_{R}G_{ss}G_{dd}+2(2E_{R}+G_{ss})G_{dd}^{2}\big{]}s_{p}^{6}$
${}-E_{R}^{3}G_{ss}^{7}\left[G_{ss}^{2}-\left(2E_{R}-G_{ss}\right)G_{dd}\right]s_{p}^{7}$
${}+E_{R}^{3}G_{ss}^{8}(G_{ss}+G_{dd})s_{p}^{8}\,,$ (72)
and
$N_{5}^{(2)}={}$ ${}16E_{R}^{6}(2E_{R}+G_{ss})^{2}(8E_{R}+G_{ss})(5E_{R}+G_{dd})G_{dd}^{2}$
${}-8E_{R}^{6}G_{ss}(2E_{R}+G_{ss})\big{[}32E_{R}^{2}G_{ss}^{2}$
${}+8E_{R}\left(15E_{R}+4G_{ss}\right)G_{ss}G_{dd}$
${}+\left(28E_{R}^{2}+32E_{R}G_{ss}+5G_{ss}^{2}\right)G_{dd}^{2}$
${}+\left(2E_{R}+G_{ss}\right)G_{dd}^{3}\big{]}s_{p}$
${}+8E_{R}^{5}G_{ss}\big{[}4E_{R}^{2}\left(4E_{R}-G_{ss}\right)G_{ss}^{3}$
${}-E_{R}\left(52E_{R}^{2}+50E_{R}G_{ss}+9G_{ss}^{2}\right)G_{ss}^{2}G_{dd}$
${}-\big{(}28E_{R}^{3}+64E_{R}^{2}G_{ss}+31E_{R}G_{ss}^{2}$
${}+3G_{ss}^{3}\big{)}G_{ss}G_{dd}^{2}$
${}+\big{(}4E_{R}^{3}-3E_{R}G_{ss}^{2}-G_{ss}^{3}\big{)}G_{dd}^{3}\big{]}s_{p}^{2}$
${}+4E_{R}^{5}G_{ss}^{2}\Big{[}2E_{R}\left(48E_{R}+13G_{ss}\right)G_{ss}^{3}$
${}+2\left(66E_{R}^{2}+45E_{R}G_{ss}+5G_{ss}^{2}\right)G_{ss}^{2}G_{dd}$
${}+\left(20E_{R}^{2}+24E_{R}G_{ss}+7G_{ss}^{2}\right)G_{ss}G_{dd}^{2}$
${}+\left(2E_{R}+G_{ss}\right)^{2}G_{dd}^{3}\Big{]}s_{p}^{3}+4E_{R}^{4}G_{ss}^{5}\big{[}4E_{R}G_{ss}^{2}$
${}+\left(26E_{R}^{2}+25E_{R}G_{ss}+3G_{ss}^{2}\right)G_{dd}$
${}+3\left(2E_{R}+G_{ss}\right)G_{dd}^{2}\big{]}s_{p}^{4}$
${}-2E_{R}^{4}G_{ss}^{5}\big{[}\left(42E_{R}+5G_{ss}\right)G_{ss}^{2}$
${}+\left(20E_{R}+11G_{ss}\right)G_{ss}G_{dd}$
${}+2\left(2E_{R}+G_{ss}\right)G_{dd}^{2}\big{]}s_{p}^{5}$
${}-2E_{R}^{3}G_{ss}^{7}[\left(6E_{R}+G_{ss}\right)G_{ss}$
${}+3\left(E_{R}+G_{ss}\right)G_{dd}]s_{p}^{6}$
${}+E_{R}^{3}G_{ss}^{8}(5G_{ss}+G_{dd})s_{p}^{7}+E_{R}^{2}G_{ss}^{10}s_{p}^{8}\,.$ (73)